### v1

# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

Abstract
This paper evaluates the economic impact of Renewable Energy Communities (RECs) on the Italian wholesale power market. Combining a bottom-up engineering approach with a short-run economic impact assessment, the study begins by mapping existing and emerging RECs in Italy. We identify key characteristics of RECs, such as average installed capacity, institutional profiles of members, types of renewable systems used, and distribution across Italy’s electricity market zones. This mapping yields representative REC configurations, which are employed within a bottom-up engineering model to generate energy injection and self-consumption profiles for different REC prosumer and producer categories (residential, public, small and medium enterprise, non-profit organization, and standalone installation), considering the different levels of solar irradiance in Italy based on latitude. These zonal results, aggregated on an hourly basis, inform the implementation of the synthetic counterfactual approach, which develops alternative scenarios (e.g., 5 GW target for REC-driven capacity set by Italian policy for 2027) to assess the impact of REC-driven injection and self-consumption on the Italian day-ahead power market. The findings suggest that REC deployment can increase equilibrium quantities during daylight in most of the time, while decreasing equilibrium quantities mostly during the cold months, as electrified heating drives greater self-consumption and offsets lower grid injections. Both positive and negative effects on equilibrium quantities suggest that REC deployment also has a potential to reduce wholesale electricity prices. Moreover, by reducing grid exchanges through higher self-consumption, REC proliferation can alleviate pressure on the distribution system.
Keywords: Energy Communities; Energy Policy; Energy Transition; Italian Electricity Market; Photovoltaic; Renewable Sources
JEL Classification: Q42, Q47, Q55, Q56
1 Introduction
The current integration of Renewable Energy Communities (RECs) in national electricity markets has made it necessary to accurately describe the behaviour of decentralized energy producers and consumers who unite into RECs. Depending on the historical path and national legislation, energy communities could act on the market as utilities do, thereby having the possibility to arbitrate between different market prices. However, RECs typically integrate a small number of users (Lupi et al., 2021). Hence, their direct interaction with wholesale markets is impossible or not allowed by a regulator. In such cases, members of RECs should be supplied by external utilities and be considered as simple users (or clients) of such utilities and/or external aggregators. Nonetheless, these users may exhibit various characteristics, acting as consumers, prosumers, or prosumagers (Sioshansi, 2019).
A comprehensive simulation model that optimizes energy sharing among households and municipal buildings within a REC is provided by Casalicchio et al. (2025), supporting REC planning through sector-coupling and economies of scale. Magni et al. (2025) investigated Italy’s transition from experimental to definitive policies for RECs, by focusing on the economic impacts implied for different RECs configurations across regions. The authors found that the current policy framework reduces regional disparities but may discourage investment in the South of Italy, despite its solar potential. Crucially, understanding the composition of prosumers (users who both generate and consume energy) within RECs (and seasonal heterogeneity) is key to investigate the role of RECs in reshaping market price signals, balancing local demand, and reducing reliance on the grid.
RECs have been deployed in many parts of the world. Europe has the largest number of RECs, with almost 4,000 initiatives and 900,000 members (Koltunov et al., 2023). European RECs deserve particular attention because of the scale of the phenomenon and their formal recognition at the EU level by the RED-II (European Union, 2018) and IEM Directives (European Union, 2019), which triggered their further growth. Outside Europe, RECs exist in the USA, Canada, Costa Rica, Australia and New Zealand, where they are mostly represented by community solar utility initiatives or historic rural cooperatives (Kolesar, 2022). In many states of the Global South such as Brazil and India, RECs are established as off-grid microgrids in areas with limited or no preexisting access to the grid (Hochstetler and Born, 2022; Thapar et al., 2017). Due to the lack of legislation and large heterogeneity of practices, RECs outside Europe remain a niche sector, thus restricting the potential for their deployment (Koltunov and De Vidovich, 2025). Therefore, the impacts of RECs on electricity markets may be negligible, and thus hard to investigate in such jurisdictions.
RECs are undoubtedly considered a unique instrument for a just energy transition. According to the EU legislation, there are two core features that define a REC:
  • RECs are the entities that should provide social, environmental and economic benefits to their members and respective local communities rather than solely for financial gain.
  • RECs should be effectively controlled by shareholders or members who are located in close proximity to the renewable energy projects owned and developed by REC’s legal entity (European Union, 2019, 2018).
There are several obvious economic and social benefits for citizens willing to join RECs and for local communities hosting them. For example, prosumers and consumers could benefit from a collective generation facility and from shared energy, while local communities usually benefit from the redistribution of profits derived from REC activities (e.g., investment in local infrastructure, nature-based solutions, energy education projects). Economic impacts of RECs can be various, ranging from effects on members and investors, on local and regional economies, and on electricity market and its stakeholders.
Table 1 depicts a taxonomy of economic impacts of RECs. The pre-Directives literature (Bauwens et al., 2016; Brummer, 2018; Brummer et al., 2017; Candelise and Ruggieri, 2017; Heras-Saizarbitoria et al., 2018; Holstenkamp and Kahla, 2016; K.Huntala, 2016; Koltunov and Bisello, 2021; Kooij et al., 2018; Magnani and Osti, 2016; Magnusson and Palm, 2019; Tricarico, 2018; Vernay and Sebi, 2020; Wirth, 2014; Walker et al., 2010; Berka and Creamer, 2018; Moroni et al., 2019) mainly explored micro-scale socio-economic impacts on REC members with a specific focus on domains of management and sociology, while rarely examining the meso [^1] Pre-Directives, there are only few economic studies with those exploring RECs impact on the regional economies using input-output models (Lantz and Tegen, 2011; Okkonen and Lehtonen, 2016; Torgerson, 2006; Phimister and Roberts, 2012; Allan et al., 2011; Bere et al., 2015; Entwistle et al., 2014) and macro effects of RECs on the economy and markets.
Table 1: Taxonomy of REC economic impacts. Source: Koltunov and De Vidovich (2025).
Category Impact on members Impact on local Impact on market stakeholders
of impact and investors and regional economies and the electricity system
Scale Micro Meso Macro
Subject One EC Multiple ECs All RECs deployed
at a country level
Objects Individual members, Local and Generators, retailers,
investors (if not a member) regional economies DSOs, aggregators,
non-member consumers,
ESCOs, technology providers
The macro impact of REC deployment on electricity systems remains an underexplored domain (Koltunov and De Vidovich, 2025). Robinson and del Guayo (2022) propose two categories: systemic impact and stakeholder impact. The first category explores the impact of RECs on the electricity systems overall, which entails multiple spillovers on the operations of various actors. Examples of such systemic impacts include changes in distribution charges by Distribution System Operators (DSOs), transmission charges by Transmission System Operators (TSOs), distribution and transmission system expansion, wholesale prices, changes in collected taxes and levies embedded in electricity tariffs, and market competitiveness (Berg et al., 2024; Frieß et al., 2025). The second category examines the impact of RECs on specific stakeholder groups, such as non-members, retailers, generators, DSOs, aggregators, and technology providers. The present article belongs to the first category, as we focus on the merit-order effect and the impact of RECs on the day-ahead (DA) power market equilibrium, which entails multiple spillovers on the operations of various stakeholders.
In contrast to the multiple benefits at the organizational and local levels, REC deployment could lead to adverse effects from an economic perspective at macro level: e.g., an increase in distribution grid charges in the presence of volumetric tariffs, which could adversely affect households non-participating in RECs. Post-Directives, several economic studies have appeared on the effect of RECs on the electricity system (Backe et al., 2022; Sarfarazi et al., 2020; Fuentes González et al., 2022; Di Silvestre et al., 2021; Boccard and Goetz, 2025). However, most post-Directive studies focus on the analysis of peer-to-peer trading, an innovative feature of REC that is yet rarely encountered in real-world applications [^2] More details on RECs classifications can be found in Rossetto et al. (2022), Kolesar (2022), Koltunov et al. (2025b) (Castellini et al., 2021; Chen and Gao, 2024; Dong and Li, 2024; Glachant and Rossetto, 2021; Hahnel et al., 2020; Hahnel and Fell, 2022; Nieto-Martin et al., 2019; Sousa et al., 2019).
By complementing existing achievements in the literature, our paper investigates the interactions of Italian RECs with the national wholesale DA market. In terms of business model archetype, Italian RECs can be classified as a ’community collective generation’. [^3] In this type of REC, generation facilities must be connected to the same voltage substation and all members should live in its proximity, while individual members (households/SMEs/public entities/non-profit organisations) maintain their own retailers that take care of their residual demand. This type of REC is the one that is mostly adopted in Italy. This type of REC includes prosumers, consumers, and producers (i.e., the collective generation facility). During periods of excess of generation, RECs supply energy to the grid, thereby increasing renewable dispatch on the supply-side of the wholesale market. At the same time, these RECs aim for higher self-consumption, especially when the selling price is lower than the retail price, thereby decreasing the aggregated load on the demand-side of the market.
For the scope of this paper, we state our research question as follows:
  • What is the impact of REC deployment on the Italian wholesale day-ahead market equilibrium?
We selected Italy as a case study for several reasons. First, Italian RECs have shown substantial growth in recent years. [^4] In 2023, there were around 50 RECs aligned with new regulation in Italy (Koltunov et al., 2023). In the beginning of 2025, there are nearly 362 RECs either in operational or design phases. Since the transposition of the EU Directives to Italian legislation, 362555Data includes RECs in both operational and design phases as of January 2025 new RECs were constructed. The primary factor behind this notable deployment is the premium tariff, granted at the end of 2023 (MASE, 2023). To date, this incentive is the only available subsidy for new photovoltaic installations in the country. Second, the Italian wholesale day-ahead power market is well-suited for our research because of its transparency regarding hourly supply and demand bids placed by market operators. Third, Italian regulation is rather innovative and can be characterized as a “virtual scheme”, where REC members (prosumers and consumers) do not physically share energy in a microgrid. [^6] In contrast, CEER (2019) reports that “Some Member States, such as France and Austria, have developed a framework for collective self-consumption, where energy can be shared within a group of customers, without requiring direct involvement of a supplier.” Finland adopted a concept of energy sharing that is typically limited to apartment blocks or single housing associations. Similar to Finland, German “collective self-consumers” concept relates largely to occupants of the same building or small groups rather than a rule allowing multi-building or multi-substation virtual sharing across a distribution grid. Instead, they retain their existing retail contracts while sharing energy virtually, for which they are remunerated at the end of the year by the GSE [^7] GSE - in Italian, Gestore del Sistema Energetico - is the energy system manager, a public company that manages incentives in the energy sector on the state’s behalf. (Schiavo et al., 2022). This innovative regulation provides an interesting case for investigating its implications. According to new rules (ARERA, 2022), REC members can benefit in four ways: a premium tariff for shared energy; valorization of avoided grid usage – calculated based on shared energy due to connection to the same primary substation; energy sales; and energy self-consumption (Blasuttigh et al., 2025). Figure 1 summarizes the benefits.
Notably, only the sales of injected energy (by prosumers and producers) and the self-consumed energy (by prosumers) affect the wholesale market, with the former impacting the supply-side and the latter the demand-side. The incentivized ‘shared energy’ is merely a virtual concept that does not directly trigger wholesale market changes. Therefore, we analyze only the operations of REC members that directly influence the market – namely, prosumers and producers – and disregard consumers. [^8] When joining a REC, consumers do not necessarily change their consumption profile in the ‘virtual scheme’ applied in Italy, therefore the wholesale market outcome does not amend. Nonetheless, a REC manager/aggregator might advice or order consumers to increase/decrease consumption in certain hours for the maximization of the common shared energy. Such advanced coordination requires smart infrastructure at the consumers’ premises and could potentially affect the wholesale market situation. However, to the authors’ best knowledge, no RECs in Italy have employed such level of coordination yet. Therefore, this aspect lies out of the scope of the current study and remains for further investigation.
Figure 1: Economic benefit to members of RECs in Italy (excl. explicit capital subsidy).
We apply an innovative mixed-model strategy in our study. In the first stage of analysis, we design a bottom-up engineering model that simulates the hourly behavior of various typical prosumers and producers within a REC. Then, we project the model’s results onto several scenarios based on the actual REC deployment status in 2024 and the policy target for 2027. The outcome of the first stage is then inserted into our synthetic modelling approach, which simulates short-run schedule for the DA market on an hourly basis, from which we derive the effect of RECs on the merit order and aggregated demand.
Hence, the novelty of our work is twofold. First, we provide a theoretical contribution to the understanding of the impact of different types of RECs on the electricity system – specifically, at the wholesale power market level –, where the changes in the equilibrium quantity and price due to REC deployment have multiple spillovers on the operations of other generators, retailers, and final consumers. Notably, heterogeneous REC membership becomes typical in many EU countries, especially after the implementation of the RED-II and IEM Directives. However, the existing literature on modeling RECs and prosumers’ behaviour mostly simulates the impacts of residential prosumers at a single geographical location, whereas our study simulates RECs as a collection of prosumers with very different consumption profiles (public, residential, small and medium enterprises, non-profit organizations) and standalone producers at various climatic locations, which closely reflects the country-scale deployment situation. Our contribution is tested empirically for the entire year 2024 in the DA Italian pool using a synthetic approach to simulate real-world competition among market agents. By undertaking such a detailed modeling exercise, our theoretical contribution may achieve higher external validity, especially for jurisdictions and markets with similar technical definition of a REC (e.g., Spain, Greece, Czech Republic), climate conditions and generation profiles (e.g., Spain, Greece).
Second, we provide a novel mixed-model approach to assess the country-scale impact of prosumers and producers aggregated into a REC on wholesale market scheduling. The first stage of our methodology begins with detailed data collection aimed at deriving available information on REC compositions, generation plant capacities, consumption profiles of different categories of prosumers, and the geographical distribution of RECs. By utilizing real-world data, we design the engineering model to obtain assumed prosumers’ generation and load profiles, which we then project to the policy-targeted deployment level (“Policy Scenario”) and other predicted deployment levels (”Half-way Scenario”, ”Business-as-usual Scenario”). In the second stage, the methodology is expanded by incorporating the previous outputs into an economic synthetic model. In a nutshell, our methodological approach can be used by policymakers and interested stakeholders to determine the actual and predicted systemic effects of REC deployment.
The remainder of the article is organized as follows. In Section 2, we explore the existing literature on systemic impacts of RECs. It then concludes with a review of studies that utilized a methodology similar to ours, i.e. the ”synthetic approach” for modeling electricity markets. Section 3 dives into the methodology employed in this paper, while Section 4 displays the major results from the analysis. Section 5 embeds the results into the academic debate, discusses their implications for policymaking and future research directions. Section 6 concludes with final remarks.
2 Literature review
Due to the recent introduction of Energy Communities, few papers address the specific mechanisms by which RECs interact with the wholesale market. Few theoretical studies based on narratives (Robinson and Arcos-Vargas, 2023; Robinson and Del Guayo, 2022; Williamson, 2022) speculate briefly on the direction of the systemic impact of RECs on wholesale quantity and prices. Moreover, we have found only four empirical studies on the matter (Backe et al., 2022; Fuentes González et al., 2022; Sarfarazi et al., 2020; Boccard and Goetz, 2025). Several empirical studies quantify the effect of individual prosumers (e.g.(Chen et al., 2023a, b; Tsybina et al., 2023)) as well as individual and aggregated prosumers (Riaz et al., 2019) on wholesale power market operations. However, quantifying the effect of prosumers on the wholesale market is not the core objective of these studies. Another limitation of these works is that the authors test their models on a small subset of actual markets (usually only a limited number of nodes/buses) over a short time scale (e.g. one week).
The work of Riaz et al. (2019) on the effect of aggregated prosumers on the wholesale market is conceptually the closest to the approach adopted in the present paper. However, apart from the methodological approach, a critical conceptual difference is that the optimization model in Riaz et al. (2019) was tested on residential prosumers, whereas our mixed model uses outputs from the operation of various categories of prosumers united in a REC – including residential, public, non-profit organizations, and small and medium enterprises (SMEs) – as well as producers. Faia et al. (2021) take an opposite approach compared to our paper. The authors explore the effect of the wholesale market on prosumers’ bills by optimizing the minimum electricity volume required for prosumers to participate in both the wholesale and retail markets, suggesting that if households are aggregated (e.g. by a REC), they can achieve significant cost reductions.
This Section reviews the relevant literature on the topic, by exploring two main areas:
  1. 1.
    Studies that analyse the effect of RECs on electricity systems;
  2. 2.
    Studies that employ a counterfactual approach for modelling the impact of specific power generation sources (e.g., RES) on the wholesale power market.
We have also reviewed studies that analyze the effect of prosumers on electricity systems and discuss their relevant findings in Section 5. However, since they do not constitute the central role in our paper, unlike studies that focus on the effect of REC deployment on the system, we omit reporting them in the main text although including them in the Appendix 7.3.
2.1 Impact of RECs on electricity systems
Several authors use a theoretical narrative-based approach to discuss the impacts of RECs (Biggar and Hesamzadeh, 2022; Del Pizzo et al., 2022; Parag and Sovacool, 2016; Robinson and Del Guayo, 2022; Di Silvestre et al., 2021). Robinson and del Guayo (2022) analyse REC features that potentially favour the system and its stakeholders, as well as those that may be harmful, using Spain as the case study. Biggar and Hesamzadeh (2022) emphasize the regulatory challenges arising from RECs and discuss various policy responses. Del Pizzo et al. (2022) examine criteria for REC sizing and geographical boundaries from a DSO’s perspective, aiming to align them with the needs of other grid users. Parag and Sovacool (2016) analyze the benefits and caveats of three categories of energy communities: peer-to-peer REC, microgrid REC, and aggregator REC. Di Silvestre et al. (2021) review REC regulations around the world and outline possible challenges and opportunities for power systems that arise from REC proliferation.
Several studies indicate systemic benefits of REC deployment. Encouraging investment in embedded generation at a larger scale increases economic efficiency (Biggar and Hesamzadeh, 2022). Locating RECs in places with under-supply of electricity reduces transmission grid costs and losses due to an increased energy supply where it is mostly needed (Robinson and Del Guayo, 2022). RECs reduce the need for an additional large-scale generation due to the merit-order effect, which in turn reduces the wholesale price and the need for additional transmission infrastructure (Robinson and Del Guayo, 2022; Backe et al., 2022). Robinson and Guayo (2022) argue that RECs can provide ancillary services if aggregated into virtual power plants, thereby replacing conventional sources which are more expensive to operate in many cases. When RECs are aggregated and utilize BESS, they could reduce peak heat load and total heat demand in highly electrified systems (Backe et al., 2022). New ESCOs and aggregators could emerge as facilitators between RECs and the grid, therefore enhancing residential and commercial energy efficiency efforts (Parag and Sovacool, 2016). Finally, multiple social benefits (Biggar and Hesamzadeh, 2022; Robinson and Del Guayo, 2022) and improved competitive pressure on other market agents (Robinson and Del Guayo, 2022) are also important systemic benefits.
At the same time, other studies point to systemic detriments of REC deployment. Biggar and Hesamzadeh (2022) and Robinson and Guayo (2022) argue that a REC may arbitrate [^9] Arbitrage effect emerges when collective generators and consumers/prosumers have distinct selling and buying prices. When they unite into an energy community, generator now can supply energy directly to members of the REC while latter remunerate the former with a payment that is higher than conventional price for selling into the grid but lower than a conventional buying price for consumers. As a result, both parties win while the regulatory effect of the price has vanished. between retail and selling prices in cases when it acts as a united entity which purchases part of its electricity externally from a single retailer. As a consequence, the effect of non-uniform tariffs (time-of-use and locational tariffs) is eliminated, which could make the system operation costlier and potentially may lead to the further expansion of the network. Biggar and Hesamzadeh (2022) indicate that retail customers do not typically face the correct incentives to use energy according to market and grid situations. When customers unite into RECs it may further amplify the impact of such inefficient tariffs. The same study (2022) argues that the reduction in network exploitation exacerbates network operator’s ‘death spiral’ effect [^10] Death spiral effect is a well-known phenomenon which describes a situation when distributed energy resources lead to prosumers paying less with volumetric distribution charges. Consequently, system operators (DSOs/TSOs) become under-paid for its previous infrastructure investment, which is usually higher than a decreased usage of a network due to self-consumption. System operators shift these costs instead onto other consumers, which in turn stimulate them to transform into prosumers as well, thereby shrinking the customer base iteratively.. Robinson and Guayo (2022) and Parag and Sovacool (2016) say that locating RECs in places with over-supply of electricity increases transmission grid costs and losses due to increased electricity injection where congestion bottlenecks are present. In addition, small capacity renewable generators are costlier for the system because they are more expensive by LCOE than large capacity renewable generators (Robinson and Del Guayo, 2022). Parag and Sovacool (2016) point that RECs might erode sensitive protections on privacy. Lastly, the same authors argue that some REC business models might be harder to deploy than others. For example, it is difficult to align the interests of peer-to-peer RECs with the interests of a broader society. Peer-to-peer business model can also entail high transaction costs. As we can see, the literature points to multiple potential harms to the system and stakeholders [^11] Since our study aims to analyze systemic impacts, we skip the literature review of impacts on the individual stakeholders in the main text. Instead, we include the review of these impacts in Appendix B. from the deployment of RECs. However, these potential harms relate to a variety of REC types and not all are necessarily applicable to the ‘community collective generation’ type deployed in Italy  [^12] Koltunov (2025a) can be consulted for an in-depth discussion on the impacts of individual REC business models on the electricity system and its stakeholders..
Fuentes Gonzalez et al. (2022) utilize a game theoretical approach with optimization modelling for a simplified three-nodal Chilean market to find the potential impact of REC deployment on a consumer surplus, nodal prices, social welfare, generation and transmission expansion, and CO2 emissions. The study reveals that the nodal price decreases from 65.77 \$/MWh to 65.4 \$/MWh when RECs are deployed in a scenario with 9.24 MW of installed capacity with 150 MW peak nodal load. Moreover, the price drops further up to 64 \$/MWh if a REC’s installed capacity gets higher, though contingent upon the disposable income of REC members. The reduction in the nodal price suggests that the effect of RECs can also be positive on the non-member consumers. An increase in social welfare can be observed when RECs are involved in energy production (14,221 \$/hour with RECs versus 14,596 \$/hour without REC) due to the avoidance of high generation costs at nodes without REC, as well as the less frequent need for transmission expansion. This leads to a reduction in the nodal price that triggers an increase in the quantity demanded from 105.79 MWh to 115.03 MWh with an energy community. Quantities demanded in their static model are simulated for a single representative hour while prices are simply derived from linear demand functions. Consequently, this approach does not take into consideration a full interplay of technologies and variability over multiple time periods that are inherent in pay-as-clear zonal wholesale auctions of the real world.
Sarfarazi et al. (2020) also use a game-theoretical approach with optimization and dynamic programming to simulate the interaction of a stylized German REC (that consists of flexible consumers, inflexible consumers, prosumers, prosumagers, and community energy storage) with its retailer and a wholesale market. The study indicates when community storage is owned by the retailer who has the profit maximization objective the scenarios with real-time purchase and feed-in selling tariffs do not decrease the quantities dispatched at the supply side of the wholesale market by RECs compared to the scenario with static tariffs when community storage decreases energy dispatched at the supply side of the wholesale market. In addition, in this scenario REC operations become aligned with the wholesale market price signals. Moreover, real-time tariffs with installed community energy storage increase profits for REC members. Therefore, this scenario leads to welfare gains both for the system and a REC. Conversely, if a retailer pursues a REC self-sufficiency maximization objective and not a profit maximization objective, REC imports less electricity from the public grid, therefore, less of grid charges are collected for the DSOs and less of levies to the general budget. Sarfarazi et al. (2020) focus primarily on the ”REC-retailer” interactions on a micro scale and do not investigate REC deployment on a macro scale, MOE and complexities of a DA wholesale market.
Boccard and Goetz (2025) test empirically the profit sharing rules for a condominium energy community and a REC in three geographical locations with diverse solar irradiation levels [^13] Weak irradiation in Hamburg (Germany), medium irradiation in Girona (Northern Spain), high - in Faro (Southern Portugal) and for different PV sizes that correspond to 25%, 50%, and 100% load coverage of households in a REC. The study finds that REC members decrease their bills in all simulated scenarios. However, the impact on the grid can be diverse. When a PV system is sized at 25% of the aggregated load of REC members, power exchanges with the grid drop from 70 MWh (no PV scenario) to 59 MWh in Northern Spain, which is a 14% decrease compared to the scenario without a PV system. This is a significant finding which indicates that a DSO could manage the deployment of RECs and generally DERs without having to raise tariffs and without compromising on grid reliability. When the PV capacity is increased to cover 50% of the aggregated load, power exchanges with the grid decrease by 7%. In contrast, in the scenario where a PV capacity covers 100% of the member load, power exchanges with a distribution grid raise up to 29%. The situation worsens in the northern irradiance region (Germany), while improves in the southern irradiance location (Portugal). The opposite pattern was revealed by the sensitivity analysis. In the scenario with heterogeneous public buildings (schools and hospitals) that install PV systems covering 50% and 100% of the demand, the power exchanges reduce at 8% and 13%, respectively, compared to no PV scenario. Conversely, the 25% PV coverage of public loads raises power exchanges with the grid at 10%. A similar situation occurs with the higher number of more heterogeneous residential loads. The authors argue that the deployment of small-scale PV systems in residential RECs could decrease operation costs for the grid, while large-scale PVs force costly upgrades within buildings and across a distribution network. Therefore, the sizing of PV systems in RECs should be carefully calibrated not only from the perspective of economic benefits for REC members but also considering systemic impacts.
Last but not least, Backe et al. (2022) utilized two large optimization models to explore the effect of the REC deployment on the additional capacity needs in a 2060 carbon-neutral Europe. The authors find that the REC deployment reduces both electricity and heating costs and lessens the need for national capacity expansion by 50-60 GW in six EU countries. However, they do not quantify the merit-order effect (MOE) or REC interaction with the DA market, openly stating that this is beyond the scope of their paper.
2.2 Synthetic approach to simulating electricity markets
While optimization models only represent in detail the behavior of a firm, equilibrium and simulation-based models represent market behavior considering competition among all participants (Ventosa et al., 2005). Equilibrium models are based on the definition of the equilibrium which is mathematically expressed in the form of a system of algebraic or differential equations, which imposes limitations on the representation of competition between participants and are frequently too hard to solve (p. 5, 2005). Simulation models are an alternative to equilibrium models when the problem is too complex to be addressed by researchers within a formal equilibrium framework (p. 6, 2005). Therefore, simulation models provide a more flexible way to address the market problem than equilibrium models, which justify its usage for our research purpose. However, a limitation of the simulation models is that they are based on assumptions that are particular to each study.
The study by Sensfuß et al. (2008) is one of the first in which the “synthetic supply” approach was applied to analyze the impact of renewable electricity generation on the spot market equilibrium. This approach involves modeling a counterfactual scenario without generation from RES and comparing it with the actual scenario where generation from RES is present. The authors find that, between 2001 and 2006, the price reduction due to the merit-order effect from renewable generation in Germany was significant, reaching its peak in 2006 at approximately €5 billion. The paper concludes that the economic benefit of RES proliferation is greater than the cost of subsidies. The net profit for consumers was €1.9 billion in 2006 alone.
A prominent topic in the simulation-based literature is the measurement of market power. For example, Ciarreta and Espinosa (2010) apply a “synthetic supply” approach in the Spanish electricity market to show that actual market prices were about 21% higher than those in a counterfactual scenario without strategic bidding by large firms, especially in 2002 and 2005. Rossetto et al. (2019) conduct a similar exercise for the Italian market between 2015 and 2018, finding that consumer surplus losses attributable to the dominant operator grew over time and were most pronounced during peak-demand months.
Beltrami et al. (2021) examine the merit-order effect in the Italian market, demonstrating that RES displace conventional generators and thereby lower overall wholesale prices. Notably, by subtracting subsidies from the environmental and economic benefits, the study still finds a 44% net welfare increase in 2018 alone, which is similar to findings from the Sensfuß and Genoese (2006) for Germany. Espinosa and Pizarro-Irizar (2018) similarly employ a synthetic supply approach to estimate the social costs and benefits of Spain’s RES subsidies over the 2002–2017 period. They show that while the subsidies initially provided net social benefits, cuts to these subsidies eventually led to reduced RES deployment, which in turn diminished the merit-order effect and offset many of the gains. Turning to electricity storage, Beltrami (2024) investigates pumped hydro storage in Italy. The ‘synthetic’ simulation reveals that the CO2 saved during discharge exceeds the CO2 generated during charging.
Synthetic approach has also been applied to hypothetical market scenarios. Ciarreta et al. (2024) explore a prospective Moroccan electricity market, comparing two counterfactual market designs to a status quo scenario without a liberalized market. They find price reductions of 48% and 43%, largely driven by more efficient dispatch and interconnection usage with Spain, and emphasize the importance of block bids in lowering final prices.
3 Methods and data
For clarity, we divided our methodological approach into two main stages. The first stage begins with mapping all Italian RECs to derive key input parameters for the subsequent engineering modeling. This model calculates both the energy injected into the grid and the energy self-consumed by various categories of prosumer load profiles, based on irradiation levels across different market zones. The model’s output is used to project REC deployment for the year 2024 (ex-post situation) and for the year 2027 (predictive situation), according to three main scenarios: best case, mid-case, and worst case. This projection step provides the input variables for the second stage of analysis.
In the second stage, we first simulate a market equilibrium with RECs that reflects the actual market situation, using Italian day-ahead market bids for the year 2024. Then, we construct a counterfactual scenario of market equilibrium without RECs for both 2024 and 2027. As the most recent annual data is available only for 2024, and the policy target is set for 2027, we assume the wholesale market conditions in 2027 to be identical to those of 2024 for simplification purposes. This assumption does not compromise our modeling objective, which is to introduce a new methodology for estimating the impact of REC deployment on wholesale market equilibrium. Our aim is not to forecast the actual impact of REC deployment in the policy-target year. Nonetheless, policy-relevant insights can still be drawn from the predicted situation. In the first stage of methodology, we used a combination of software applications: MS Excel (mapping), Matlab (engineering modeling), R (projection). In the second stage, R was utilized as the main modeling software. Each stage is described in more detail below.
3.1 First stage: Mapping, Engineering modeling, Scenarios and Projections.
3.1.1 Comprehensive Mapping of Italian RECs
To accurately model the actual and potential impact of RECs, we began by collecting data on all operational and planned RECs in Italy. [^14] The complete data on the mapping of Italian RECs are available in Supplementary materials. Due to the absence of a comprehensive dataset from a single source, we compiled our database from multiple sources, specifically:
  • the data portal of GSE [^15] The state-owned renewable energy agency
  • the annual reports of Legambiente [^16] A national environmental non-governmental organisation, i.e. NGO.
  • the data portal NeXt [^17] NeXt — the civil society network comprising the majority of Italian third-sector and public bodies working in the REC field ESG
  • the data portal Sinergie Condivise [^18] Developed by a banking foundation ‘Fondazione Compagnia di San Paolo‘ in collaboration with regional and municipal governments and universities in the Piedmont Region
  • publicly available business plans of individual RECs
  • academic publications
  • websites of REC developers
  • websites of news agencies, regional and municipal authorities, and individual initiatives.
Our final database contains 34 variables for 362 RECs in Italy, of which 184 are in the operational phase and 178 are in the design phase as of 30 January 2025. However, for the purposes of this study, we used only the following 20 variables:
  • electricity market zone
  • number of primary substations to which the REC is connected
  • capacity (installed or planned), kW
  • PV capacity (installed or planned), kWp
  • capacity per substation, kW
  • self-consumption level, %
  • battery availability and capacity, kWh
  • number of members and installed capacity per category of prosumer/consumer/producer
  • building types per category of prosumer
  • capacity share of public prosumers in total REC capacity, %
  • capacity share of residential prosumers in total REC capacity, %
  • capacity share of SME prosumers in total REC capacity, %
  • capacity share of NPO prosumers in total REC capacity, %
  • capacity share of PV producers in total REC capacity, %
  • number of rooftop installations by public prosumers
  • number of rooftop installations by residential prosumers
  • number of rooftop installations by SME prosumers
  • number of rooftop installations by NPO prosumers
  • number of standalone PV installations
  • installed renewable energy technology.
Zhu et al. (2025) also built the database of new Italian RECs for the purpose of their study. In addition to the diverse scope of variables collected (due to different research objectives), another major difference lies in the development status of the RECs that were included in both databases. Our database includes both operational RECs and those that were in the project design stage of development as of January 2025. Inclusion of RECs in the project design stage allowed us to have a more holistic perspective on the future trends, therefore, to construct robust scenarios grounded in the larger sample of the real-world data, which is discussed in further detail in Section 3.1.3. The detailed comparison of two databases can be found in Table 3.1.1.
Comparison between the databases Variable Our database Zhu et al. (2025) Total number of RECs, from which: 362 212 Operational RECs 184 212 Design-phase RECs 178 0 Location 362 212 Population (of municipality) 85 Climatic zone 85 Market zone 362 Generator’s capacity 300 212 Type of technology installed/planned 323 85 Number of membersa 283 Category of membersb 219 Share of installed PV capacity, by category of prosumers 271 Technical indicatorsc 147 BESS availability 13 7 EVs and/or e-charging availability 38 3 Self-consumption level 82 Economic indicatorsd 105 Number of prosumer buildings by type 158 55 Number of consumer buildings not counted 41 Investment source 252 CAPEX costs 106 Legal form 180 REC builders and promoters 204 Last update January 2025 February 2025 Data sources GSE, Legambiente, RSE, Sinergie, Condivise, NeXt ESG, business plans, academic publications, websites of builders, other websites GSE, Legambiente, RSE, academic publications
  • a
    Members include technical users (prosumers, consumers, producers), and non-user members.
  • b
    Categories of members include: public, residential, SME, NPO.
  • c
    Technical indicators collected from business plans include energy generated, injected, and shared as well as self-consumption level.
  • d
    Economic indicators collected from business plans include revenues from energy sales and energy sharing, savings from energy self-consumed.
From our database, we obtained input parameters for the subsequent methodological stages. The first input parameter can be calculated via the following equation:
$p_{a,one}^{P_{as}}=\frac{p_{a,total}^{P_{as}}}{N_{r}^{P_{as}}/N_{p}^{P_{as}}}$ (1)
From our database, we derived input parameters for the subsequent methodological stages. The first input parameter can be calculated via the following equation:
$P_{p,z}^{PV,avg,one}=\frac{P_{p,z}^{PV,avg,total}}{N_{p}^{r}}$ (2)
where
  • $p$ denotes a category of prosumer/producer.
  • $z$ denotes an electricity market zone.
  • $Z$ is the set of electricity market zones, where $Z\in\{\text{North},\text{Central North},\text{Central South},\text{Calabria},\\ \text{South},\text{Sicily},\text{Sardinia}\}$. [^19] From now, we refer to market zones through their codes NORD, CNORD, CSUD, CALA, SUD, SICI, SARD.
  • $P_{p,z}^{PV,avg,one}$ is the average PV capacity per one prosumer/producer of category $p$ in electricity market zone $z$.
  • oneone denotes one prosumer/producer
  • $\mathcal{P}$ is a set of prosumer/producer categories, where $\mathcal{P}\in\{\text{Public},\text{Residential},\text{SME},\text{NPO},\\ \text{Standalone producing installation}\}$.
  • $P_{p,z}^{PV,avg,total}$ is the total average PV capacity of all prosumers/producers of category $p$ within an individual REC in market zone $z$.
  • $N_{p}^{r}$ is the average number of rooftop or standalone installations for prosumers/producers of category $p$ across the entire country [^20] Due to the small sample sizes for the number of rooftop/standalone installations across individual zones in our database, we decided to estimate the average number of rooftop/standalone installations across the entire country, that is, $N_{p}^{r}$. However, if the sample size allows, the methodology should instead employ the averages across individual zones, that is, $N_{p,z}^{r}$ .
In turn, the array of $P_{p,z}^{PV,avg,total}$ values was derived via the following equation:
$P_{p,z}^{PV,avg,total}=\frac{1}{n_{p,z}}\sum_{n=1}^{n_{p,z}}P_{p,z}^{PV,ind}\left(\frac{Sh_{p,z}^{ind}}{100}\right)$ (3)
where
  • $P_{p,z}^{PV,ind}$ is the PV capacity of all prosumers/producers of category $p$ within an individual REC located in market zone $z$.
  • indind refers to an individual REC.
  • $Sh_{p,z}^{ind}$ is the capacity share of prosumers/producers of category $p$ within an individual REC in zone $z$.
  • $n_{p,z}$ is the number of RECs that include at least one prosumer/producer of category $p$ in market zone $z$.
Similarly, the self-consumption level of 49.1% was identified as the average across all RECs for which such data were available. Consequently, our engineering model is based on a self-consumption range of 45%–50%–55%, where the central value corresponds to the real-world situation, the lower bound reflects a slightly more pessimistic scenario, and the upper bound represents a more optimistic scenario that reflects the potential deployment of battery energy storage systems (BESS). However, only 3.6% of the RECs currently report having plans to install BESS. Although BESS would enable both a higher self-consumption rate and increased remuneration for shared energy, long payback periods inhibit their adoption.
Another input parameter is the ’most common building type’, which is essential to select realistic prosumer load profiles in the engineering model. For example, public prosumers may be represented by the load profile of a school or a sports facility, while SME prosumers may be modelled using the load profiles of commercial buildings or hotels. Clearly, different load profiles yield different results. Therefore, identifying the most common building type is critical for producing outputs that closely reflect reality without introducing excessive complexity by modelling all possible building types. This parameter was derived from our database through manual counting of building types for different categories of prosumers.
3.1.2 Bottom-up Engineering Modelling of RECs
Figure 2: Graphical representation of the prosumer energy modeling framework used to simulate hourly energy flows for different prosumer/producer categories. The flowchart illustrates the input data (PV production and load profiles), the modeling process for each prosumer/producer category (residential, school, commercial, office, standalone), and the generation of annual hourly datasets for energy production, self-consumption, and grid injection across multiple market zones and self-consumption scenarios. These outputs are then used as inputs for the scenario-based projection and economic simulations in the day-ahead electricity market.
Modeling these entities at a fine-grained level allows a detailed assessment of how different categories of users interact with their renewable energy systems over time. This insight is crucial for calculating quantities such as self-consumed energy and injected energy into the grid, which are useful for determining the change in energy volumes and thus for estimating economic market variables. An overview of the proposed bottom-up engineering methodology is shown in Fig. 2, which is explained in the following paragraphs.
To develop our behavioral energy model of prosumers, we designed a modeling scheme that represents five prosumer/producer categories, namely residential, schools, commercial, office, and standalone PV systems. Except for the standalone PV systems (which act purely as generators without any demand), all user categories are equipped with both PV production and electric load, connected to the national power grid.
From real-world data collected across operating RECs in Italy, as described in the previous Section 3.1.1, we estimated the average PV nominal capacity installed for each prosumer category within different RECs. These values were then used to scale hourly-based annual PV production profiles, obtained from location-specific PV yield profiles for seven representative cities covering the main Italian market zones: Milan for NORD, Florence for CNORD, Rome for CSUD, Brindisi for SUD, Catanzaro for CALA, Palermo for SICI, and Cagliari for SARD.
Consumption patterns for each prosumer were modeled using nPro (Wirtz, 2023), a load profiling tool capable of generating hourly demand curves based on user-specific behavior. To explore different energy scenarios, the load profiles were scaled to achieve three predefined self-consumption levels: 45%, 50%, and 55%.
This approach resulted in the generation of annual hourly-based time series (8,760 values) for each prosumer/producer category and each scenario, capturing PV generation, electricity consumption, self-consumed energy and energy injected into the grid. With seven market zones and three levels of self-consumption considered, the final output of the engineering model consists of twenty-one structured datasets. Each dataset includes the energy behavior of the four prosumer categories and the standalone PV systems, whose entire production is fed into the grid. These profiles are then used as input for the subsequent projection step, which is described in Section 3.1.3.
Photovoltaic Generation Profiles
To characterize the hourly production of PV systems across Italy, we made use of the PVGIS platform developed by the European Commission’s Joint Research Centre (PVGIS, 2022). For each of the seven electricity market zones we selected a reference city as defined above, and extracted location-specific PV yield data. Rather than relying on generic assumptions, we configured each PV system in PVGIS under optimal operating conditions. This meant that, for each site, the tilt angle of the PV modules was set to the value maximizing annual output, the azimuth orientation was chosen to face due south, which typically yields the best yearly performance in Italy and the PV technology considered was crystalline silicon, the most widespread module type. Also, system losses, such as reflecting factors, inverter efficiency, temperature effects, cable losses, dust, and shading, were included using an average value of 15% (Ogliari et al., 2023).
For each location, the tool provides an hourly-based annual time series representing the energy yield per installed kilowatt-peak. These values were then scaled by the average installed PV capacity for each prosumer/producer category $P_{p,z}^{PV,avg,one}$, based on data from existing RECs, as in (4):
$E^{PV,one}_{p,z}(t)=Y^{PV,one}_{p,z}(t)\cdot P_{p,z}^{PV,avg,one}$ (4)
where
  • $Y^{PV,one}_{p,z}(t)$ [kWh/kWp] is the hourly-based annual yield profile of the reference city from PVGIS of category p in electricity market zone z for one prosumer/producer
  • $E^{PV,one}_{p,z}(t)$ [kWh] is the hourly-based annual profile of energy produced by the PV system of category p in electricity market zone z for one prosumer/producer
  • t is the time index $\in\{1,2,...,8760\}.$
By using PVGIS in this way, we were able to obtain consistent, reproducible photovoltaic production profiles that reflect regional climatic differences and common technical configurations, without introducing excessive complexity into the modeling process. It’s worth noting that, while real-world data from monitored PV systems might offer greater accuracy in principle, such datasets are often fragmented, inconsistent, or not openly available across all regions. Simulated data from PVGIS, on the other hand, ensures full spatial coverage and comparability, while still being grounded in satellite-based irradiance data and validated performance models.
Load Consumption Profiles
To represent electricity consumption behavior for each prosumer category within the REC framework, we generated hourly demand profiles using nPro (Wirtz, 2023), a profiling tool that provides synthetic yet behaviorally-informed load curves. The tool produces time series based on statistical models of daily and seasonal usage patterns for different consumer categories (residential, schools, commercial, office, and many others) under typical operational conditions  [^21] All prosumers are assumed to own electrical loads related to heating, cooling, and general electricity demand, except for residential prosumers, for which only cooling has been considered in order to model summer air-conditioning. For all loads, the default values provided by nPro have been retained, meaning that no additional calibration of consumption profiles was performed beyond the standard dataset assumptions. . However, while these base profiles are useful in capturing the temporal distribution of consumption during the year, they do not by default reflect a specific relationship with local PV generation. In particular, no predefined level of self-consumption (i.e., the portion of PV energy that is immediately consumed by the user) can be assumed unless demand and generation are explicitly aligned.
To introduce variability in this key parameter and explore its influence on community-level energy flows, we implemented a simple but effective adjustment strategy. The idea was to scale the demand profiles vertically to increase their overall magnitude, without altering their temporal pattern. This approach changes the extent to which demand and generation tend to coincide and allows us to control the self-consumption ratio, a value that is typically chosen when designing PV systems, in straightforward way.
Specifically, for each user category p in electricity market zone z, the original load profile generated by nPro was multiplied by a constant factor which was iteratively tuned until the ratio of self-consumed energy to total PV production matched a given target. We considered three such targets (45%, 50% and 55%) chosen to represent a plausible range of self-consumed energy. In this setting, energy not consumed at the time of generation is immediately injected into the grid, as no energy storage system is supposed to be available.
However, to account for the likely evolution of REC configurations in the coming years, we also explored additional scenarios featuring higher self-consumption levels. These extended cases are intended to emulate the effect of widespread adoption of energy storage systems, such as residential or commercial batteries. By shifting consumption toward daylight hours or enabling the deferred use of solar generation, energy storage reduces the amount of surplus energy injected into the grid, thus increasing the local use of renewable electricity. Although modelling of storage systems is not explicitly included in this framework, these scenarios of increased self-consumption provide an indication of their aggregate effect on energy flows. This allows an exploratory assessment of how the progressive penetration of storage technologies may alter the energy balance and the economic impact of RECs on the market.
Calculation of Self-Consumed and Exported Energy
Once both the hourly PV generation and the electricity demand profiles were established for each prosumer/producer category and scenario, the next step involved calculating the two profiles that define the interaction between local generation and the grid: self-consumed energy and surplus energy injected into the main grid. These quantities were derived directly from the hourly time series of PV production $E^{PV,one}_{p,z}(t)$ and electrical load $E^{L,one}_{p,z}(t)$ of category p in electricity market zone z computed as explained in the previous section. In particular, the self-consumed energy $E^{self,one}_{p,z}(t)$ of category p in electricity market zone z for one prosumer/producer at each hour is computed as in (5):
$E^{self,one}_{p,z}(t)=min(E^{PV,one}_{p,z}(t),E^{L,one}_{p,z}(t))$ (5)
This corresponds to the portion of the PV production that is immediately used to meet on-site demand. When the load exceeds the available PV power, all PV generation is consumed locally. Conversely, if generation exceeds demand, only part of it is self-consumed. In this case, the surplus energy exported to the grid $E^{exp,one}_{p,z}(t)$ of category p in electricity market zone z is calculated as in (6):
$E^{exp,one}_{p,z}(t)=max(E^{PV,one}_{p,z}(t)-E^{L,one}_{p,z}(t),0)$ (6)
In other words, the exported energy is the excess generation that is not used locally and is therefore injected into the public distribution network. These expressions were applied element-wise over the full annual time series, resulting in two additional vectors of hourly values for each prosumer/producer and scenario.
In the end, for each prosumer category, the combined simulation produces four time series per user: hourly PV production, electricity consumption, self-consumed energy and energy fed into the grid. Instead, for standalone PV producers, the energy injected into the grid is equal to the PV energy produced, hour-by-hour.
Figure 3: Hourly energy profiles for Milan (NORD) over a full year for four prosumer user categories: residential, school, commercial, and office. Each subplot shows PV generation (yellow line), electricity consumption (blue line), and energy injected into the grid (green line).
Figure 3 shows an example of the time profiles explained in the previous paragraphs for Milan (NORD). In particular, the annual hourly energy profiles for four different categories of prosumer users are presented. In each subplot, three key variables are shown over the full time horizon of one year. The yellow curve represents the electricity generated by the PV system, while the blue line indicates the hourly electricity consumption associated with the user. The green line, plotted below the horizontal axis, corresponds to the surplus energy that is not self-consumed and is instead injected into the grid. To make the graphs easier to read, the hourly profile of self-consumed energy is intentionally not shown.
The energy behavior varies significantly across prosumer categories. Residential prosumers exhibit relatively low and stable generation throughout the year, with a load profile that allows only limited self-consumption, especially during daytime hours. School buildings show a highly intermittent demand, with pronounced reductions in summer months and strong peaks during the winter, reflecting the academic calendar. Commercial prosumers show a more consistent and uniform demand profile, enabling a better alignment between PV production and consumption. Also, office buildings exhibit similar patterns, although with lower overall demand and more visible weekday-weekend variability. In all cases, the amount of feed-in is visibly related to the time mismatch between photovoltaic production and load demand.
3.1.3 Scenarios and Projection of REC Deployment
The core objective of the simulation framework is to evaluate the market impact of RECs under a range of realistic deployment pathways. The calibration of the scenarios is primarily based on the fulfillment - or lack thereof - of Italy’s policy target of 5 GW of installed REC-linked RES capacity by year 2027 as well as on the current REC deployment levels (367.514 MW of PV capacity in operational and design status; 29.646 MW in operational status). Three main narrative trajectories are hereby defined:
  • Policy Scenario: this scenario assumes the full achievement of the 5 GW target by 2027, reflecting an optimistic rollout of RECs in terms of both regulatory support and investment mobilization.
  • Half-way (HW) Scenario: we assume the steady proliferation of the 2024 deployed capacity, with 0.368 GW in operational and design status, reaching 1.47 GW by 2027.
  • Business-as-usual (BU) Scenario: this scenario assumes no significant acceleration in REC deployment beyond the current trend, with 29.646 MW in operational status in 2024. For the 2027 projection, we again assume the steady proliferation of the 2024 operational capacity, reaching 0.119 GW.
Table 2: List of simulated scenario assumptions. sc = self consumption.
No. Scenario name Scenario code Year % of sc assumed GW of REC installed capacity Policy target achieving
1 Policy sc45.2027 2027 45% 5 Yes
2 Policy sc50.2027 2027 50% 5 Yes
3 Policy sc55.2027 2027 55% 5 Yes
4 Half-way sc45.HW.2027 2027 45% 1.47 No
5 Half-way sc50.HW.2027 2027 50% 1.47 No
6 Half-way sc55.HW.2027 2027 55% 1.47 No
7 Business-as-usual sc45.BU.2027 2027 45% 0.119 No
8 Business-as-usual sc50.BU.2027 2027 50% 0.119 No
9 Business-as-usual sc55.BU.2027 2027 55% 0.119 No
10 Mixed scenario 1 sc_mix1.2027 2027 Public: 50% Residential: 45% SME: 55% NPO: 50% Standalone: no sc No
11 Mixed scenario 2 sc_mix2.2027 2027 Public: 55% Residential: 55% SME: 55% NPO: 55% Standalone: no sc Yes
Within each of these trajectories, scenarios are further differentiated based on assumed self-consumption rates, defined as the share of generated renewable energy consumed by prosumers within a REC rather than injected into the grid. This parameter is a critical driver in the modelling framework, as it reflects the efficiency and “virtuosity” of RECs in optimizing on-site energy use - thus affecting both demand-side and supply-side market dynamics. Additionally, two “mixed scenarios” are included to reflect more granular assumptions based on prosumer-category segmentation, as identified through our empirical mapping of REC profiles. These scenarios apply differentiated self-consumption rates by prosumer category (e.g., residential, public, SME, NPO), thereby enhancing the realism and policy relevance of the simulated outcomes. A summary of all scenarios is provided in Table 2.
The inputs for the scenario-based projections were provided by the engineering model that yielded two key variables related to individual prosumers/producers within a REC: annual energy self-consumed, $E_{self}(t)$, and the energy injected into the grid, $E_{export}(t)$, by typical prosumer categories and standalone PV systems. To represent the additional dimensionality of these variables, namely prosumer category and market zone as well as their individual nature, we denote them as $E_{p,z}^{\mathrm{self.REC}}(t)$ and $E_{p,z}^{\mathrm{export.ind}}(t)$ further in text. These values for self-consumed and injected energy were projected for both the ex-post (2024) and predicted (2027) deployment scenarios (see Table 2), using a set of parameters derived from the collected database.
The first projection parameter, “Zonal Shares of RECs,” represents the percentage distribution of RECs across electricity market zones:
$Sh_{z}^{total}$ $=\frac{100}{n}\sum_{n_{z}=1}^{n_{z}}n_{ind}$ (7)
where
  • $Sh_{z}^{total}$ denotes the zonal share of RECs,
  • $n_{ind}$ is an individual renewable energy community,
  • $n_{z}$ is the total number of RECs in each zone $z$, where $z\in Z$,
  • $n$ is the total number of RECs in Italy, $n=362$.
The “average capacity share of all REC prosumers/producers” is the second projection parameter. It is calculated using the following equation:
$Sh_{p,z}^{avg}$ $=\frac{1}{n_{p,z}}\sum_{n_{p,z}=1}^{n_{p,z}}Sh_{p,z}^{ind}$ (8)
where
  • $Sh_{p,z}^{avg}$ is the average capacity share of all REC prosumers/producers of category $p$ in market zone $z$.
  • $Sh_{p,z}^{ind}$ is the capacity share of prosumers/producers of category $p$ within an individual REC in zone $z$.
  • $n_{p,z}$ is the number of RECs where a prosumer category $p$ is present in zone $z$
The total REC deployed PV capacity per prosumer category $p$ and zone $z$ was calculated for each scenario using the equation:
$P_{p,z}^{PV,total}$ $=\frac{P_{scen}^{PV,total}\times Sh_{z}^{total}\times Sh_{p,z}^{avg}}{100}$ (9)
where
  • $P_{p,z}^{PV,total}$ is the total REC deployed PV capacity per prosumer category $p$ in market zone $z$,
  • $P_{scen}^{PV,total}$ is the scenario-based total REC deployed PV capacity.
The number of PV plants per prosumer category $p$ and market zone $z$ can be calculated using:
$Plant_{p,z}$ $=\frac{P_{p,z}^{PV,total}}{P_{p,z}^{PV,avg,one}}$ (10)
where
  • $P_{p,z}^{PV,avg,one}$} is the average PV capacity installed by one REC per prosumer category $p$ in market zone $z$.
By taking the outputs of the engineering model - $E_{p,z}^{\mathrm{exp,one}}(t)$ and $E_{p,z}^{\mathrm{self,one}}(t)$ - and knowing $Plant_{p,z}$, we can derive the projected energy injected into the grid and the projected energy self-consumed by different categories of prosumers/producers:
$E_{p,z}^{exp}(t)$ $=E_{p,z}^{\mathrm{exp,one}}(t)\times Plant_{p,z}$ (11)
$E_{p,z}^{self}(t)$ $=E_{p,z}^{\mathrm{self,one}}(t)\times Plant_{p,z}$ (12)
where
  • $t$ is a specific hour in a year, $t\in T$ and $T=\{1,2,\ldots,8760\}$,
  • $E_{p,z}^{exp}(t)$ is the projected energy injected into the grid during hour $t$ by all photovoltaic plants of prosumer category $p$ in market zone $z$,
  • $E_{p,z}^{\mathrm{exp,one}}(t)$ is the energy injected into the grid during hour $t$ by one photovoltaic plant of prosumer category $p$ in market zone $z$,
  • $E_{p,z}^{self}(t)$ is the projected energy self-consumed during hour $t$ by prosumer category $p$ in market zone $z$,
  • $E_{p,z}^{\mathrm{self,one}}(t)$ is the energy self-consumed during hour $t$ by one prosumer of category $p$ in market zone $z$.
Finally, we derive two arrays of projected variables: (i) energy injected into the grid by RECs, and (ii) energy self-consumed by RECs, using the following equations:
$E_{z}^{exp}(t)$ $=\sum_{p\in\mathcal{P}}E_{p,z}^{exp}(t)$ (13)
$E_{z}^{self}(t)$ $=\sum_{p\in\mathcal{P}}E_{p,z}^{self}(t)$ (14)
where
  • $E_{z}^{exp}(t)$ is the projected energy injected into the grid by all RECs during hour $t$ in market zone $z$,
  • $E_{z}^{self}(t)$ is the projected energy self-consumed by all RECs during hour $t$ in market zone $z$.
3.2 Second stage: Economic modeling
To assess the short-run economic impact of RECs on the Italian wholesale power market, we apply an empirical hour-by-hour counterfactual simulation similarly to the work by Beltrami et al. (2021), further extended by Beltrami (2024). This methodology builds on publicly available data from Gestore del Mercato Elettrico – GME – namely, “Offerte Pubbliche” (or public offers) – to reconstruct the merit-order demand and supply curves for each of the Italian electricity market zones and each hourly settlement period. From now onwards, in order to maintain the focus on the synthetic approach, we employ a simplified notation which is consistent with the formal used in Subsection 3.1. For instance, $\Delta Q_{\text{REC,d}}$ corresponds to $E_{z}^{self}(t)$, while $\Delta Q_{\text{REC,s}}$ corresponds to $E_{z}^{exp}(t)$ in Eq. (13) and (14). The reference to time and zone is omitted. The strategy holds for all zones and settlement periods. Our counterfactual simulation aims to model the hypothetical configuration of the day-ahead market in the absence of REC-driven generation and self-consumption.
This is a formal description of the theoretical approach we adopted. We consider an electricity market where, given a price level $Pr$, the demand curve is denoted by $D_{actual}(Pr)$ and the supply curve by $S_{actual}(Pr)$. These functions are derived from the aggregation of individual bids and offers submitted in a uniform-price day-ahead auction. The market equilibrium is determined by the intersection of these curves at the market-clearing price $Pr_{actual}$, such that:
$D(Pr_{actual})=S(Pr_{actual})=Q_{actual}$
where $Q_{actual}$ is the corresponding equilibrium quantity. In this setting, synthetic supply and demand curves can be constructed to represent a competitive benchmark, using observed bidding behavior or reconstructed marginal costs across all operators. These curves serve as the baseline for analyzing the effect of RECs.
As stated above, RECs typically self-consume a share of their electricity production. This self-consumed electricity is not visible in the observed market demand. In a counterfactual scenario where RECs are absent, this hidden demand would need to be satisfied by the wholesale market.
Let $\Delta Q_{\text{REC,d}}$ denote the total quantity of REC self-consumption. To account for this, the demand curve is adjusted by shifting it horizontally to the right:
$D_{synt}(Pr)=D(Pr)+\Delta Q_{\text{REC,d}}$
On the supply side, RECs may inject electricity into the market through surplus generation. Removing RECs implies that this contribution is also removed from the market supply. Let $\Delta Q_{\text{REC,s}}$ represent the total quantity of REC-generated electricity that would have been offered to the market. The adjusted supply curve is thus defined as:
$S_{synt}(Pr)=S(Pr)-\Delta Q_{\text{REC,s}}$
This corresponds to a horizontal shift of the supply curve to the left.
The counterfactual equilibrium price and quantity in the absence of RECs are denoted by $Pr_{synt}$ and $Q_{synt}$, respectively. These are determined by the intersection of the adjusted demand and supply curves:
$D_{synt}(Pr_{synt})=S_{synt}(Pr_{synt})$
Substituting the shifted functions, the equilibrium condition becomes:
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}=S(Pr_{synt})-\Delta Q_{\text{REC,s}}$
This can be rearranged as:
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}+\Delta Q_{\text{REC,s}}=S(Pr_{synt})$
This framework captures the dual effect of RECs on market equilibrium (see Figure 4). The removal of self-consumption by RECs increases observed demand, while the removal of injection by RECs reduces available supply, thus producing a rightward shift in the demand curve and a leftward shift in the supply curve. As a result, the counterfactual equilibrium price $Pr_{synt}$ is expected to be higher than the baseline price $Pr_{actual}$. The change in equilibrium quantity, $Q_{synt}-Q_{actual}$, depends on the relative elasticities of supply and demand. This formalization enables a quantitative evaluation of the role of RECs in lowering prices and reducing market dependency in wholesale electricity auctions.
Specifically, we modify the observed merit-order curves as follows:
  • Supply curve shift. In hours where RECs inject renewable electricity into the grid, we assume that this volume would not be available under the counterfactual scenario. Accordingly, the supply curve is shifted leftward to reflect the reduction in total market supply, primarily from RES. This leads to a counterfactual configuration in which electricity prices would be higher, ceteris paribus.
  • Demand curve shift. In hours where RECs self-consume a portion of their generation, the equivalent electricity demand is effectively removed from the market. In the counterfactual scenario, where self-consumption does not occur, we expand the demand curve rightward (upward shift in the merit-order framework), capturing the higher residual market demand that would otherwise materialize.
Figure 4: Graphical representation of the market equilibrium taking into account the effect of RECs. The curves $D_{\text{actual}}$ and $S_{\text{actual}}$ represent the actual demand and supply with RECs, leading to a market equilibrium at price $Pr_{\text{actual}}$ and quantity $Q_{\text{actual}}$. The dashed curves $D_{\text{synt}}$ and $S_{\text{synt}}$ refer to synthetic demand and supply, resulting in an alternative equilibrium at price $Pr_{\text{synt}}$ and quantity $Q_{\text{synt}}$.
This dual (contemporaneous) adjustment is performed hourly and for each market zone, generating synthetic demand and supply curves representing a no-REC baseline condition. The resulting counterfactual price and quantity outcomes are then compared with the observed market outcomes to estimate the impact of REC operations.
We acknowledge that this simulation is based on several assumptions. First, the counterfactual is computed under a ceteris paribus condition, assuming all other market dynamics unchanged. [^22] This assumption is particularly relevant when comparing the effects of RECs on demand and supply across the different scenarios, and becomes evident mostly in interpreting the results from the Policy Scenario compared to BU and HW. In our setting, we implicitly assume that the installed capacity of other RES technologies (beyond RECs) remains fixed. This does not reflect the actual dynamics of the Italian power system, where RES capacity additions are currently progressing at sustained rates as well as does not reflect the demand growth. Nevertheless, this choice is consistent with our primary goal, i.e. to isolate the specific marginal contribution of REC deployment, without counfouding their individual effect with broader renewable expansion trends. Second, each market zone is treated as a “closed system”, abstracting from inter-zonal electricity flows. While this assumption may limit economic and policy implications, it remains defensible for two main reasons: (1) RES-generated power is typically prioritized in dispatch due to the merit-order principle, making its theoretical removal analytically legitimate; (2) the current scale of REC operations in Italy is still limited, rendering its impact negligible in terms of strategic bidding by large market players deeply affecting the market clearing dynamics. Overall, this synthetic control approach allows us to isolate and quantify the localized, short-term effects of REC deployment on market outcomes such as zonal clearing prices and traded quantities, thus offering a robust empirical basis for evaluating REC-driven welfare gains.
4 Results
4.1 Descriptive statistics of REC configurations in Italy
Table 3 displays the completeness of the available data. The information on the market zone and the type of installed technology is the most complete. Crucial data on the nominal capacity of power plants is available for 82.9% of RECs. The estimation of many parameters required data on the category of members, which is available for 74.9% of RECs. The typology of prosumer buildings is available for only 43.7% of RECs. In turn, the self-consumption level—from which the range of 45–50–55 scenarios was assumed—is available for only 22.7% of RECs. Similarly, only 3.6% of RECs plan to install BESS.
Table 3: Available data on RECs used for the study
Available Information Number of RECs
Total number of RECs 362
Information on the electricity market zone where a REC is located 362
Information on technology installed in a REC 323
Information on generator’s capacity installed in a REC 300
Information on category of members
(consumer, prosumer [public, residential, SME, NPO], producer)
271
   from which:
    Category of members and member number 219
    Only category of members without number 52
Building types where a prosumer’s facility is installed
(e.g., school, hospital, condominium, town hall, church, etc.)
158
Self-consumption level of REC prosumers 82
Battery availability and/or capacity in a REC 13
Figure 5 shows that most RECs are situated in the NORD market zone (63.8%) and the Central-South zone (19.1%). In contrast, RECs are scarcely present in the rest of Italy. This deployment pattern may be associated with the general distribution of economic activity across the country. Since the second half of the 20th century, Northern Italy and the regions surrounding Rome have exhibited high levels of industrial and entrepreneurial activity. Consequently, the greater availability of expertise and financial resources for REC establishment has supported their rapid proliferation in these two market zones.
Figure 5: Zonal shares of RECs, $Sh_{z}^{\text{total}}$.
Figure 6: Distribution of RECs across prosumer/producer categories
In Figure 6232323Figure 6 contains information not for all 362 RECs from our database but for RECs with the available information., we observe that the most common prosumer/producer category is public prosumers—typically represented by municipal buildings. Standalone systems are reported for 61 RECs, while rooftop installations owned by SMEs and residential prosumers are reported for 42 and 36 RECs, respectively. In contrast, only 22 RECs report involving NPO prosumers. The higher number of RECs with public prosumers may be explained by the fact that municipalities are, by far, the most common promoters of RECs in Italy, for whom targeted capital and regional grants are also more accessible. In addition, public buildings consume most of their load during daylight hours, making them ideal candidates for maximizing the state incentive on shared energy. Accordingly, standalone plants also allow for the maximization of profits due to their greater average installed capacity, especially when balanced with the daytime loads of public, SME, and NPO prosumers and consumers. Although residential prosumers have not yet actively participated with their rooftop systems, many residential consumers [^24] REC consuming members are not the focus of this study. Therefore, even though data on REC consumers is available, we do not report it here. are members of RECs in Italy [^25] The elaborate discussion on equity concerns and private citizen participation in Italian RECs can be found in Koltunov (2025b)..
Figure 7: Average capacity shares of all categories of prosumers/producers across market zones, $Sh^{a}_{p,z}$
Figure 7 shows the average capacity shares of all categories of prosumers/producers across market zones. Photovoltaic installations of all categories are present only in three zones: NORD, CSUD, and SUD. The CSUD and NORD market zones have the most proportionate prosumer/producer capacity distributions. This situation is possibly associated with the high number of RECs located in these zones (see Fig. 8). Moreover, the strong economic activity in these regions may contribute to a greater diversity of stakeholders who possess the financial, administrative, technical, legal, and social engagement resources and skills necessary to participate in RECs (Koltunov, 2025b). In contrast, a more disproportionate distribution of capacity across prosumer/producer categories - where public bodies dominate - can be observed in the market zones of Southern Italy (SUD, SARD, SICI, CALA). This differentiation into REC prosumers’ participation is also supported by the qualitative findings of Musolino et al. (2023). Importantly, the absence of installed plants for certain member categories does not automatically exclude them from REC membership. For example, all member categories may also participate in RECs as consumers without owning rooftop systems.
Figure 8: Average PV capacity per one Prosumer/Producer, $P^{a.one}_{p,z}$.
Note: The number after the zone name indicates the sample size (REC observations).
Figure 9: Building/installation types by prosumer/producer categories.
Note 1: The numbers after building types indicate the number of individual buildings.
Note 2: “Others” in the Public category include police stations, autodromes, waste management companies, cemeteries, public utility facilities, etc. “Others” in the NPO category include social canteens, sports and educational centers, and social farms.
Figure 8 demonstrates the “Average PV capacity per one Prosumer/Producer“, $P^{a.one}_{p,z}$, which is one of the two real-world input parameters for the engineering model. [^26] Table in Appendix A reports values for two constituent parameters, $P^{a.total}_{p,z}$ and $N^{r}_{z}$, that are used to derive $P^{a.one}_{p,z}$ The average capacity of a single photovoltaic plant varies across market zones. For public and SME prosumers, the average capacity in the NORD and Central zones (CNORD, CSUD) is significantly higher than in the Southern zones (CALA, SUD, SICI). [^27] Only 5 observations were used to estimate the average SME capacity in SARD. Therefore, 334.5 kW of average capacity installed by SMEs in SARD is based on a limited sample and should be interpreted with caution. Similarly, the very large capacity of public prosumers in CNORD zone, 285.5 kw, may be related to the small observational sample. Again, this pattern may be associated with easier access to financing in Northern and Central Italy. Another notable observation is that NPOs have significantly smaller installed capacities on average compared to SMEs and public prosumers. Typically, non-profit organizations have more limited access to private financing than SMEs, while capital grant subsidies are reserved exclusively for small municipalities (Koltunov, 2025b). Finally, zero values are mostly observed in the Southern zones, possibly due to the small number of REC observations.
In Figure 9, we observe the types of buildings where PV installations have been constructed by prosumer category. This information was used to identify the “most common building type.” For public prosumers, data on specific building types is unavailable for most RECs; consequently, the most common identified building type is a school. For SMEs, it is a commercial building, typically a supermarket or shopping mall. Hotels and industrial buildings are also relatively common in the SME category. The type of almost all residential buildings is unspecified (935 buildings), followed by private detached houses (14 buildings). Offices and churches are the most common building types for the NPO category. Finally, 65 standalone photovoltaic plants have been reported, while non-photovoltaic technologies have been used in a much smaller number of facilities.
We also analyzed the average number of buildings participating as prosumers in a single REC, $N^{r}_{p}$. On average, 4.4 public buildings with rooftop plants participate in a REC. Only 2.3 SME buildings and the same number of NPO buildings participate as prosumers in a REC. In contrast, approximately 17.7 residential houses participate with rooftop systems in a REC. However, as shown in Figure 8, residential prosumers have, on average, much smaller generating capacities. The average number of standalone producers per REC is 1.7.
4.2 Results from the synthetic approach
This subsection displays the core results of the paper, presenting main outputs at weekly, monthly, and quarterly resolutions. Results are reported exclusively for NORD and CSUD zones. [^28] Our simulation algorithm did not retrieve valid outcomes for the remaining five zones, mainly due to data discontinuities and the low diffusion of RECs in such zones.
The analysis focuses on the comparative dynamics between observed market outcomes and their synthetic counterfactuals, highlighting the extent to which REC-driven injections and self-consumption might influence zonal market equilibria in terms of traded volumes and prices.
To improve the readability of plots representing the percentage impact of RECs on equilibrium quantities in the DA market, we applied a smoothing procedure using a centered simple moving average. Specifically, for each hour of the day, we replaced the original values with the average computed over a 7-hour symmetric window, using the rollmean function from the zoo package in R. The smoothing was applied selectively, only to positive values, while non-positive entries were left unchanged. This method reduces short-term fluctuations and highlights general trends across the different scenarios and seasons. This procedure has been applied to all Figures from 14 to 23.
4.2.1 Results for main scenarios
[…]
In Figure 10, the notable difference between the magnitude of the equilibrium effects in winter (smaller spike density) and other seasons (greater spike density) can be observed. A combination of the low solar irradiance while high-self consumption in winter induces this trend. First, lower solar irradiation in winter lead to fewer quantities of energy to be injected into the grid by RECs. Second, self-consumption rate by RECs is higher during winter. Figure 11 illustrates daily self-consumption levels for all prosumer categories in different months. All categories self-consume on average more energy in January than in April, July or October. As a result, the RECs’ self-consumption effect on the demand side of the market is relatively bigger than the RECs’ injection effect on the supply side of the market in winter than in other seasons. In addition, low solar irradiance during winter aggregates the effect of offset of energy injection with energy self-consumption. In our engineering model, we added electrified heating and cooling to the load profiles of public (schools), SME (commercial), and NPO (offices) categories [^31] However, residential prosumers own just electrified cooling systems in our model and not heating systems, similar to the actual status quo in Italian households.. Therefore, in Figure 11, we observe that heating needs increase self-consumption during winter, mostly, for only three, albeit REC dominating (Figure 7), categories of prosumers.
Figure 10: Hourly impact on equilibrium quantities by RECs in the North market zone.
Figure 11: Daily self-consumption rates for all prosumer categories in 12 months of the year.
[…]
By focusing on the NORD zone and delving into seasonal variations, Figure 13 displays the average profile of actual and counterfactual market quantities by hourly settlement period for January and April 2024. Under the HW scenario assumption, RECs deliver a net positive effect on actual market quantities only for central hours of the day in January (left panel). In contrast, during a typical spring month (April), the positive effect on market quantities is already visible since 6am, lasting until 7.30pm, in line with the fading sunlight (right panel). Thus, the scale of the impact of RECs is, as expected, stronger in April, because of the larger penetration of RES from RECs and their more effective displacement of expensive thermal generation due to the merit-order effect.
[…]
We further investigate the latter finding, by analyzing the percentage relative difference between real and counterfactual hourly quantity profiles - stemming from the outcomes of Figure 13 -, and applying the smoothing procedure reported at the beginning of this Section. Intuitively, a positive percentage relative difference indicates that actual market quantities outweigh counterfactual synthetic volumes, while a negative percentage relative difference would signal the opposite effect.
Yet, Figure 14 shows that both the HW (sc45.HW.2027) and the BU (sc45.BU.2027) scenarios display a positive relative percentage effect of RECs on actual market quantities. Conversely, the Policy scenario (sc45.2027) falls under the zero line for several settlement periods - especially during peak hours -, thus suggesting that RECs have the potential to reduce traded market volumes in the Italian DA power market under large levels of deployment. Nevertheless, the magnitude of such effect has a strong seasonal dependency. Indeed, the largest magnitude of relative reduction in DA equilibrium quantities occurs in January, as a result of the aforementioned ”REC Winter Effect”. [^32] This entails a larger energy self-consumption by RECs relatively to their energy injection into the grid.
Figure 14: Percentage relative difference of average hourly impact on quantities from RECs in the NORD. The outcomes are disentangled by settlement period, and displayed by month for each designed scenario.
[…]
Figure 15: Percentage relative difference between actual and counterfactual average hourly quantities in the NORD for each month, by hour. Scenario: sc45.BU.2027.
Figure 16 shows comparable results for the HW scenario: in July, the impact of RECs on market quantities emerges earlier (4–6 a.m.), mirrors the April effect until 8 a.m., then weakens during the day, and resurfaces around 7 p.m. due to extended sunlight - similarly to the BU scenario.
As for the Policy scenario (Figure 17 ), the impact of RECs in April would oscillate around the zero-line resembling ”zig-zags” until midday, thus witnessing a balanced effect between lower market demand (due to larger self-consumption by RECs) and higher electricity supply by RECs. Instead, the impact of RECs would be more pronounced during the rest of the day. Interestingly, the effect of RECs in July is the opposite, indicating a prevailing effect of energy injected by RECs until 1 p.m., to then switch back to the dominance of the effect of reduced market demand resembling ”zig-zags”, which could be explained by the strong need of power consumption for space cooling during summer. Indeed, we notice that the green line constantly falls before dropping below zero after 2 pm. We observe that the effect of reduced market demand (explicitly evident within the Policy scenario) mimics the hourly self-consumption rates of different prosumers’ categories. For example, in July, self-consumption rates in the afternoon (1pm - 7pm) for all prosumers’ categories are higher than the morning self-consumption rates (6am - 12pm). [^33] Public prosumers raise self-consumption rates on average at 2.3%, SME prosumers at 2.7%, residential prosumers at 18.5%, and NPO prosumers at 1.1%. In contrast, in April, self-consumption rates are, on average, higher in the morning than in the afternoon. We assume that the specific composition of prosumers in RECs (Figure 7) might also affect the impact’s dynamics. However, this inquiry requires deeper investigation which is left for future research.
Figure 16: Percentage relative difference between actual and counterfactual average hourly quantities in the NORD for each month, by hour. Scenario: sc45.HW.2027.
Figure 17: Percentage relative difference between actual and counterfactual average hourly quantities in the NORD for each month, by hour. Scenario: sc45.2027.
Lastly, we break down results by showcasing weekdays and weekends effects (Figures 18-23).
[…]
Figure 18: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekdays. Scenario: sc45.BU.2027.
Figure 19: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekends. Scenario: sc45.BU.2027.
Figure 20: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekdays. Scenario: sc45.HW.2027.
Figure 21: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekends. Scenario: sc45.HW.2027.
Figure 22: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekdays. Scenario: sc45.2027.
Figure 23: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekends. Scenario: sc45.2027.
These latter findings carry important policy implications, particularly in light of the growing electricity demand driven by system-wide electrification and the potential role of RECs to reduce system stress. Yet, we argue that, in the scenario where RECs are widely deployed (Policy scenario), the system would recognize the role of RECs in reducing wholesale power demand through increased self-consumption. This would not only shield members of RECs from market price volatility, but also reduce the need for costly grid infrastructure investments, thereby facilitating a broader integration of RECs into the energy system, while consumers’ autonomy would be enhanced.
4.2.2 Sensitivity analysis: mixed scenarios
We hereby report a sensitivity analysis, by relaxing the assumption of the homogeneous self-consumption rate for all types of prosumers. Yet, we explore the impact of the two ”mixed” scenarios as indicated in Table 2, hence modeling heterogeneous self-consumption rates for prosumers’ categories.
[…]
Figure 24: Profile of average hourly quantities in NORD for both actual and counterfactual scenarios. Period: January 2024. Comparison between mixed scenarios: sc_\_mix1.2027 vs. sc_\_mix2.2027.
Figure 25: Hourly percentage impact on equilibrium quantities by RECs: comparison between mixed scenarios: sc_\_mix1.2027 vs sc_\_mix2.2027 for NORD. Period: January 2024.
5 Discussion, policy implications, and future research directions.
Italy allocated €5.7 billion to support REC deployment: €3.5 billion for premium tariffs (financed via a levy on electricity consumption) and €2.2 billion from the National Recovery and Resilience Fund to cover up to 40% of CAPEX in small municipalities. These figures highlight the strategic importance of assessing the effectiveness and system-wide value of RECs. Our paper delves into such issue, clarifying how trajectories of RECs deployment affect the equilibrium of the Italian DA power market. This Section presents an array of implications arising from our investigation and compare it with findings from other studies.
Firstly, our results show that the impact of RECs on power market outcomes is focused primarily in the NORD and CSUD market zones, with current negligible or absent effects in the other five physical market zones. This finding is closely related to the current distribution of REC projects, which are predominantly located in such two areas (see Fig. 8). Relatively slow and complex authorization and acceptance processes, institutional burdens, and limited technical support in other areas prevent broader deployment, confirming the need for targeted support policies to unlock REC potential across all regions. The similar finding has been reported by Zhu et al. (2025) and Musolino et al. (2023).
Secondly, the temporal dimension of REC impacts reveals strong seasonal and hourly patterns, driven by the interaction between solar production profiles and the assumed self-consumption rates. Our simulations for BU and HW scenarios indicate that certain early morning and evening hours - especially in spring and summer months - exhibit a relatively significant increase in market quantities, thereby highlighting the potential for RECs to better fulfill peak demand pressures in a decarbonizing system. Similarly, although including BESS, Riaz et al. (2019) modeled a summer week in the Australian wholesale market. They investigated the aggregate effect of a large number of prosumers on the load profile and found that increased prosumer participation with BESS flattens demand profiles, enhancing voltage stability and reducing the need for gas peaking plants. However, the exception in their study is the scenario with low demand and excess RES generation. It leads to increased aggregate demand due to battery charging, thus decreasing the stability margins of the system. Boccard and Goetz (2025) argue that excessive RES generation from REC prosumers entails increased power exchanges with the grid, although they do not specify in which direction.
[…]
Fourth, our results for the policy scenario indicate that RECs, by enabling higher self-consumption, could also reduce the need for costly investments in grid infrastructure. This finding aligns with the conclusions by Fuentes González et al. (2022), who modeled the effect of energy communities on transmission grid costs in Chile and found substantial system-wide savings. Backe et al. (2022) came to similar conclusions regarding the expansion of projected transmission capacity of the grid in six EU countries by 2060. Studies by Boccard and Goetz (2025) and Sarfarazi et. al. (2020) reveal that RECs could reduce pressure on the distribution grid too. Nonetheless, this effect is contingent on the size of a PV system and REC member composition in the former case and pricing strategies and community BESS operation in the latter case. While our study is not explicitly focused on grid cost optimization, the observed market impacts also suggest that RECs may alleviate pressure on the grid by decreasing energy demand from the external grid.
Fifth, nuanced PV sizing will be essential to fully realize the potential of REC deployment not only to reduce stress on the grid but also wholesale market prices. Boccard and Goertz (2025) found that the different sizing of PV systems in RECs could induce opposite impacts on the distribution system. For example, RECs with homogeneous profiles of residential prosumers, which install PV systems that cover 25% - 50% of their load, reduce power exchanges with the grid, whereas those who install PV systems that cover 50%–100% of their load increase power exchanges with the grid due to the mismatch between energy injecting and self-consuming. The latter situation could create reverse flows and thus force additional costs for the DSOs that eventually will be shifted onto the non-members of RECs. In turn, RECs with higher number of residential prosumers with heterogeneous profiles and RECs with public prosumers, which install PV systems that cover 25% of their load, increase power exchanges with the grid, while those who install PV systems that cover 50% - 100% of their load reduce power exchanges with the grid. This drastically opposite impacts on power exchanges hint at importance of modeling not only diverse PV system sizes but also diverse categories of prosumers, as we do. In our simulations, we used real-world averaged PV sizes of REC prosumers and producers in Italy (Figure 8) differentiated by categories and electricity zones. Similarly to Boccard and Goertz (2025), our future research could adopt sensitivity tests based on differentiating PV sizes, which could suggest insights related to the REC PV sizing and changes to the wholesale equilibrium. This could enrich the discussions about policies for the REC deployment in a system-efficient welfare-enhancing manner.
Finally, our work contributes to the broader policy debate on energy independence and consumer resilience in a time of extreme market volatility. The 2022 energy crisis and ongoing geopolitical tensions have reaffirmed the importance of local, decentralized energy systems. By simulating multiple realistic REC uptake scenarios, we are able to outline the range of possible benefits, while also acknowledging their limits. Even in our most optimistic scenario (sc_\_mix2.2027), the net reduction in wholesale market quantities remains moderate, suggesting that RECs are a necessary but not sufficient condition for deep decarbonization, thus requiring other structural measures and investments.
6 Conclusions
This paper fills a relevant gap in the literature, by combining bottom-up engineering estimations, an extensive review of REC deployment in the Italian setting, and an economic synthetic (counterfactual) approach based on real micro-data from market outcomes. By relying on this data-driven approach, our study delivers a novel methodological framework to assess the impacts of RECs on the electricity market, with an application to the Italian DA market.
Our results show that REC deployment generates non-negligible effects in specific zones of the market (notably North and Central-South), while its overall system-wide impact remains quite moderate. Seasonal and hourly patterns confirm that self-consumption is a key driver of economic benefits from REC deployment at macro scale. Specifically, the estimated impacts on market volumes lie within a narrow range, from a minimum of -0.19% (January, Policy Scenario) to a maximum of +1.16% (April, BU Scenario). While both BU and HW scenarios yield consistently positive but limited net effects (up to +1.16% and +0.89%, respectively), the Policy scenario reveals mixed evidence: the impacts remain close to zero in most represented months, but turn slightly negative in January, suggesting that higher REC self-consumption can, under certain conditions, reduce the volumes available in the wholesale market.
Indeed, by extending the analysis to seasonal and intra-day variations, our results confirm that the magnitude of REC impacts varies substantially across months and hours. Under the BU scenario, April emerges as the most responsive period, as market volumes with RECs surpass volumes without RECs by up to 3% during weekdays, while July exhibits a steadier impact close to 1% during peak hours. Conversely, cold months display more muted effects, with January showing negative deviations in the policy scenario due to the dominance of self-consumption patterns. Main factors behind this trend are the lower solar irradiation and the greater heating needs supplied by electricity in colder months. Lastly, when increasing the self-consumption rates for all categories of prosumers (from 45% to 55%) in line with Italian policy targets for RECs (sc_\_mix2.2027), the results yield an overall average reduction of about -0.3% on the equilibrium volumes of the Italian DA market, which is on 0.11% greater than the average reduction of -0.19% in the 45% self-consumption scenario. This finding indicates that the increase in self-consumption rates for RECs deepens the reduction in equilibrium quantities.
This latter observation suggests that by including BESS in energy communities, even greater diminishing of wholesale equilibrium quantities can be attained. The discharge of the community-owned BESS to cover the needs of REC members may entail the great systemic benefits during peak hours. An expensive and polluting thermal generation could be offset by a clean and social energy. Yet, as discussed by many authors (Soini et al., 2020; Schick et al., 2022; Chen et al., 2023b), a particular modality of a BESS system (e.g., a community-owned BESS or individual BESS for each prosumer within a REC, etc.) would eventually decide the scale of this beneficial systemic impact, although could even become adverse under certain circumstances (Sarfarazi et al., 2020).
Standalone systems are included in our REC modeling corresponding to the actual deployment status in Italy. The reducing effect of RECs’ self-consumption on market quantities also suggests that both the volumes of energy injected into the grid from renewable standalone plants and volumes from prosumers’ plants within a REC could be offset by the self-consumption of REC prosumers at the distribution system where a REC is located. This finding contributes to the debate on the effects of the renewable-dominated system with high-electrified energy demand (incl. heating, cooling and transport) on wholesale market outcomes, supporting Böttger & Härtel (2022), Backe et al. (2022) and Riaz et al. (2019) previous results, in which the authors state that the broad diffusion of distributed energy resources can reduce and smoothen aggregate demand, consequently, decreasing wholesale market prices.
This paper is not exempt from limitations. First, the calibration of scenarios relies on publicly available data on REC uptake and profiles, standard self-consumption rates, and the announced policy targets by 2027. Second, behavioral aspects and demand-side management responses are not fully captured, although they may play a decisive role in the future and in dynamic adjustments in demand profiles, mirroring peculiar weather conditions and real-time modifications. Third, the projected impact of REC deployment is evaluated under the assumption that the capacities of other RES technologies remain fixed. This allows us to isolate the marginal contribution of the REC deployment to market equilibrium, although the actual system impacts could differ if broader RES additions were accounted for. Fourth, we disregard consumers who participate in RECs, although theoretically they might amend its consumption profiles to maximize the incentive for ”shared energy” when joining RECs. However, as stated in Section 1, this degree of coordination has not yet been implemented in the Italian energy communities, moreover, at a substantial scale. Despite such limitations, our methodology has a wide replication potential. The synthetic counterfactual approach can be adapted to other EU (and non-EU) electricity markets to assess the role of RECs (or other distributed generation sources) on day-ahead market equilibrium. This would allow for a broader comparative insight into how local RECs can reshape aggregate electricity demand and supply, grid requirements, and market resilience.
Finally, our findings provide timely inputs for energy and environmental policy-making. The diffusion of RECs could foster energy independence and price stability, particularly in current times of geopolitical uncertainty and high energy costs, as highlighted during the 2022 energy crisis. However, their legacy will depend on complementary measures: faster permitting procedures, harmonized governance frameworks, stronger support for storage and demand-side flexibility, and access to financing resources. Italian government extended a deadline that grants €2.2 billion in subsidies for new installations within RECs to six additional months, with a possible extension (Camera dei deputati, 2025). This optimistic signal suggests that actual REC deployment in Italy by December 2027 will not only match our ”Half-Way” scenario but also make the 5 GW ”Policy scenario” more attainable.
Supplementary material
Our database of Italian RECs is available at the following GitHub link:
https://github.com/maksym-koltunov/energy-communities-Italy.git
Acknowledgements
The authors are deeply grateful to Prof. Alessandro Massi Pavan from the Department of Engineering and Architecture and to the Interdepartmental Center for Energy, Environment and Transport ”Giacomo Ciamician”, University of Trieste, for the great organizational and networking support during this study.
7 Appendices
7.1 Appendix A
Table 4: Glossary.
Acronym Description
BESS Battery Energy Storage System
CO2 Carbon dioxide
DER Distributed Energy Resources
DSO Distribution System Operator
EU European Union
GW Gigawatt
GSE Gestore Servizi Energetici
IEM Internal Electricity Market Directive
MW Megawatt
MWh Megawatt per hour
LCOE Levelized Cost Of Energy
MOE Merit order effect
NGO Non-governmental Organisation
NPO Non-profit Organisation
PV Photovoltaic
REC Renewable Energy Community
RED-II Second Renewable Energy Directive
RES Renewable Energy Sources
SME Small and medium-sized enterprises
TSO Transmission System Operator
USA United States of America
7.2 Appendix B
Table 5: Economic impact of RECs deployment on other market stakeholders
Specific impacts of RECs on market stakeholders Direction of impact Reference
On Generators
RECs deployment reduces the need for additional large-scale generation due to a merit-order effect (MOE) of renewables therefore reducing conventional producers’ surplus. Detriment Robinson & Guayo (2022), Backe et al. (2022)
RECs operating BESS could displace fossil fuel generators at a faster scale than large renewable generators without BESS due to MOE at peak demand periods, therefore reducing conventional producers’ surplus. Detriment (generators) / Benefit (system) Robinson & Guayo (2022)
RECs operating BESS reduce RES curtailment due to self-consumption at peak production congested times which allows to decrease supply of a distributed generation when grid needs it most (in turn, large renewable generators do not need to be curtailed thus preserving revenues). Benefit Backe et al. (2022)
On Retailers
Aggregator RECs operating BESS achieve cost reduction by directly participating in a wholesale market bypassing retailers thus shrinking their potential revenues. Detriment Faia et al. (2021)
When a REC acts as an intermediary between consumers and retailer, the latter lose ability to diversify risks by differentiating over tariffs to different consumer groups. Detriment Biggar & Hesamzadeh (2022)
Diminished sales from REC members elicit tariffs inflating for remaining customers, which in turn trigger latter to become more self-sufficient too. Situation leads to a financial “death spiral” of retailers. Detriment Parag & Sovacool (2016), Sarfarazi et al. (2020)
If a community energy storage is owned by a retailer, then the real-time pricing optimized based on behavior of REC members could yield profits for the retailer while not increasing costs for any type of REC members and delivering profits for flexible REC members. However, real time tariffs need implementation of EMS unavailable at scale in many countries. Benefit Sarfarazi et al. (2020)
On DSOs
When a REC acts as a united entity supplied only by a single external retailer, effect of a non-simultaneitya of contracted capacity is eliminated, which in turn reduces overall payment to DSOsb. Benefit Biggar & Hesamzadeh (2022), Robinson & Guayo (2022)
Electrotechnical criterion for defining perimeter of connection based on connection to same HV/MV substation (appr. 10000 PODs) and geographical criterion based on zip-code (appr. 1000 PODs) simplify DSO interaction with RECs. Benefit Del Pizzo et al. (2022)
Electrotechnical criterion of defining perimeter of connection based on connection to same MV/LV substation (appr. 70 PODs) and geographical criterion based on municipality belonging (appr. 10 to 850000 PODs) impedes DSO interaction with numerous or heterogeneous RECs. Detriment Del Pizzo et al. (2022)
Presence of RECs in some markets can cause operability issues and grid disruption due to a more complicated control and management schemes. Detriment Parag & Sovacool (2016)
Possibility to automatically detect and respond to actual and emerging grid problems through aggregated RECs (similar to VPPs), that may increase system’s resilience and decrease renewable energy oversupply concerns. Benefit Parag & Sovacool (2016)
On non-member consumers
DSOs can incur revenue losses due to decreased volumetric (decreased electricity purchasing by RECs) and/or fixed (non-simultaneity effect) network payments as well as possible subsidized exemption of RECs from the network payments. These revenue losses would typically be shifted onto non-member consumers. Detriment Biggar & Hesamzadeh (2022), Del Pizzo et al. (2022), Robinson & Guayo (2022), Sarfarazi et al. (2020)
Take-up of renewable energy supplied by RECs and all the subsequent MOE decrease energy prices, thereby non-members can greatly benefit. Benefit Biggar & Hesamzadeh (2022)
If RECs obtain implicit subsidy from a government, the costs are usually cross subsidized to non-members through retail bills. Detriment Robinson & Guayo (2022)
  • a
    Example: if 10 consumers each contract a 10-kW capacity-based network charge, the total contracted capacity would be 100 kW. However, due to non-simultaneity, regulation typically requires that the system only meets a combined demand of 75 kW, effectively overcharging consumers by 25 kW, which covers fixed network costs. When consumers form a REC, they contract only for 75 kW, benefiting from the non-simultaneity effect themselves and avoiding the 25-kW surcharge (Robinson and Del Guayo, 2022).
  • b
    This impact can backfire to non-member consumers because DSOs, as regulated monopolies, would typically shift the reduced revenues to non-member consumer bills.
7.3 Appendix C
Impact of prosumers on electricity systems
Ventosa et al. (2005) provide an overview of the model categories used for electricity markets analysis: optimization, equilibrium, and simulation models. We found several studies that analyze the systemic impact of prosumers. While two studies are theoretical review works that provide an extensive narrative of benefits and challenges (Robinson and Arcos-Vargas, 2023; Simshauser et al., 2023), the remaining studies rely on optimization models.
Simshauser et al. (2023) illustrate the case of Queensland state, Australia, which has the highest PV rooftop adoption rate in the world, while Robinson and Arcos-Vargas (2023) present both positive and negative effects of prosumer proliferation, focusing on Spain. Both studies outline many implications of prosumer proliferation, which often coincide with REC impacts. Some of the most pronounced implications include an adverse effect on conventional generators and challenges faced by DSOs due to a voltage rises (e.g., damage to customers’ electrical appliances). Another drawback is that retailers rapidly lose market share as prosumers penetration increases. Studies also point to an ambiguous impact on non-prosumers, who on one hand experience rising bills due to cross-subsidies and increased distribution charges, and on the other hand benefit from the decrease in the energy component of their retail bill as the result of the MOE and reductions in fuel costs. For instance, new PV installations in Spain resulted in a reduction of the wholesale electricity price of 0.01 euro for every 25 MWp installed, which, when aggregated across the overall market in 2021, led to user savings of more than 100,000 euros per year (Robinson and Arcos-Vargas, 2023). In Australia, another positive impact on non-prosumers was significantly lower installation costs due to the substantial growth of PV installation companies. In addition, DERs can lead to an increased need for ancillary services, which in turn reduces the price of these services due to greater market liquidity. Robinson and Arcos-Vargas (2023) argue that when prosumers and producers are aggregated (similar to RECs), costs of the distribution grid can drop because aggregated agents can provide flexibility services to the network — something that is almost impossible with disorganized individual installations. The proliferation of individual prosumers can decrease ohmic losses in the distribution grid, but only up to a certain level; beyond that, significant reverse flows occur, increasing losses again. Nevertheless, losses in such situations remain lower than before any DERs were deployed (2023, p. 135). An ambiguous impact, as underscored by Robinson and Arcos-Vargas (2023), occurs in terms of the security of supply when DERs do not utilize storage and are not aggregated. In this scenario, network costs can increase if a system is planned from an N-1 deterministic perspective; however, if it is planned from a probabilistic perspective, prosumers could improve security of supply even without storage or aggregation.
The first group of empirical studies on the systemic impact of prosumers attempts to quantify the associated challenges. For example, Schick and Hufendieck (Schick and Hufendiek, 2023) investigate the distributional spatial effect of the German feed-in-tariff during the period 2000-2021. Aggregated across Germany, the feed-in-tariff led to a cost shift of more than 500 million euros onto traditional consumer households. In 2021, maximization of self-consumption accounted for approximately half of this total effect. Tsybina et al. (2023) explore strategic behaviour of prosumers (exercising market power) and their response to the allocation of network losses—either to demand-side or the supply-side—as well as the impact of net metering policies. The authors determine that prosumers sell more electricity when losses are allocated to the demand base, whereas when losses are allocated to the supply base, prosumers sell less electricity. Another key observation is that lower wholesale equilibrium prices occur when network losses are allocated to the demand side due to two main factors. First, incorporating losses into the retail price (demand side) keeps selling prices higher and incentivizes prosumers to inject electricity into the grid. Second, higher retail prices encourage both consumers and prosumers to adjust their consumption patterns, leading to a more efficient use of energy, thereby reducing peak demand and the need for expensive peaking plants. Chen et al. (2023b) compare net-metering, net-billing, and benchmark policies [^34] Benchmark policy in this study assumes prosumers selling at a wholesale equilibrium price and buying at a retail price, therefore same tariffs are implied for prosumers as for other agents., examining their differential effects on various aspects of system welfare. The authors find that social surplus under the benchmark policy is significantly higher than under the other policies, making the benchmark policy the most welfare-enhancing. Under net metering, the transmission tariff would be 33% higher compared to the benchmark case, whereas the transmission tariff under net billing is closer to the benchmark case, making net billing the second-best solution. Another key finding is that wholesale social surplus (excluding prosumer surplus) deteriorates when prosumers saturate the node due to a greater number of consumers converting to prosumers—an indication of the “death spiral” effect. Chen et al. (2023a) refine the optimization model used in their previous study to investigate the impact of prosumers on transmission charges and social surpluses under the benchmark case alone. The authors assume different deployment levels of aggregated prosumers and analyse scenarios of perfect and imperfect competition, with the latter allowing prosumers to exhibit market power. The study reveals that wholesale prices decline under both scenarios due to increased renewable dispatch and reduced demand by prosumers. However, under imperfect competition — where prosumers strategically maximize their individual economic optimum — a significant increase in transmission charge occurs at all levels of deployment, particularly in scenarios with a high saturation of prosumers at nodes. This, in turn, reduces overall welfare.
The second group of empirical studies examines solutions to mitigate network costs shifting onto non-prosumers. Schick et al. (2021) demonstrate that network allocation schemes based on peak-coincident network capacity utilization can more effectively incentivize distribution network-oriented behaviour while ensuring a fairer distribution of financial burden between prosuming and non-prosuming households compared to volumetric network charges. A subsequent study by the same authors (2022) finds that higher self-consumption, when operated at least partially in a grid-beneficial manner (e.g., coupled with storage capable of providing flexibility service), can enhance RES integration and reduce CO2 emissions while avoiding cost shifting onto consumers. This finding suggests that dispersed prosumers could contribute more effectively to the grid if coordinated through a REC. On the other hand, when prosumers focus solely on maximizing their individual economic optimum— without considering system economic optimum (e.g., when storage operation is entirely inflexible)—RES integration could decrease, leading to a substantial rise in system costs and CO2 emissions (Schick et al., 2022). These findings align with those of Chen et al. (2023b).
The third group of studies explores future scenarios characterized by a high penetration of renewables, including DERs. Böttger and Härtel (2022) investigate hypothetical German power day-ahead market in 2050, assuming the deployment of carbon-neutral electricity/heat/transport systems. Importantly, their study considers the role of various novel demand-electrification technologies, which contribute to both supply and demand. The authors find that variable RES market values can be stabilised by power demand from diverse electrification applications, including flexible storage, power-to-gas, and power-to-heat (heat pumps). Consequently, a fully renewable future does not necessarily imply the “cannibalization effect” and highly volatile wholesale prices. Soini et al. (2020) investigate the impact of prosumers’ BESS on power supply costs in the Swiss electricity market for 2030, comparing it to the status quo in 2015. Their findings indicate that when BESS operation is optimized from a power system perspective—through time-of-use tariffs, grid charging, power exchange minimization, and households’ aggregation via RECs—substantial cost savings can be achieved. These savings primarily result from the reduced generation requirements and the substitution of pumped-hydro storage with more efficient BESS. Conversely, when fully independent households optimize their self-consumption, costs increase. This outcome is aligned with the findings of both Schick et al. (2022) and Chen et al. (2023b). Finally, Riaz et al. (2019) analyze the effect of large-scale prosumer aggregation (including BESS) on wholesale demand positions and load profiles, with a specific focus on loadability [^35] Loadability - maximum amount of a load that a system can support before it collapses. and voltage stability. Their study reveals that the increased prosumer-BESS participation smooths demand profiles, enhances loadability and voltage stability, and reduces gas power plants utilization—-thus lowering wholesale electricity prices. However, in scenarios of low demand and excess RES generation, these benefits do not occur. In contrast, a higher RES penetration without BESS leads to reverse flows and a reduction in reactive power support capability, ultimately lowering system stability margins.