### v2

# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

Abstract
This paper evaluates the economic impact of Renewable Energy Communities (RECs) on the Italian wholesale power market. Combining a bottom-up engineering approach with a short-run economic impact assessment, the study begins by mapping existing and emerging RECs in Italy. We identify key characteristics of RECs, such as average installed capacity, institutional profiles of members, types of renewable systems used, and distribuiton across Italy’s electricity market zones. This mapping yields representative REC configurations, which are employed within a bottom-up engineering model to generate energy injection and self-consumption profiles for different REC prosumer and producer categories (residential, public, small and medium enterprise, non-profit organization, and standalone installation), considering the different levels of solar irradiance in Italy based on latitude. These zonal results, aggregated on an hourly basis, inform the implementation of the synthetic counterfactual approach, which develops alternative scenarios (e.g., 5 GW target for REC-driven capacity set by Italian policy for 2027) to assess the impact of REC-driven injection and self-consumption on the Italian day-ahead power market. The findings suggest that the REC deployment can increase equilibrium quantities during daylight over most of the time, while decreasing equilibrium quantities mostly during the cold months. The differing load profiles of institutional REC members, compared to residential members, drive most of the decreasing effect. Both positive and negative effects on equilibrium quantities suggest that REC deployment also has a potential to reduce wholesale electricity prices. Moreover, when reducing grid exchanges, REC proliferation can alleviate pressure on the transmission system.
Keywords: Energy Communities; Energy Policy; Energy Transition; Italian Electricity Market; Photovoltaic; Renewable Sources
JEL Classification: Q42, Q47, Q55, Q56
Word count: 13118
1 Introduction
The current integration of Renewable Energy Communities (RECs) in national electricity systems has made it necessary to accurately describe its market consequences. Depending on the historical path and national legislation, energy communities could act on the market as utilities do, thereby having the opportunity to arbitrate between different market prices. However, RECs typically integrate a small number of users (Lupi et al., 2021). Hence, their direct interaction with wholesale markets is impossible or not allowed by a regulator. In such cases, members of RECs should be supplied by external utilities and be considered as simple users (or clients) of such utilities and/or external aggregators. Nevertheless, these users may exhibit various characteristics, acting as consumers, prosumers, or prosumagers (Sioshansi, 2019), thus implicitly affecting electricity markets. Understanding the composition and behaviour of prosumers within RECs—including their seasonal and geographical heterogeneity—is essential for assessing how these configurations reshape wholesale market price signals and reduce user dependence on the grid. Such insights are critical for designing policies that generate benefits not only for REC participants but also for the broader energy system.
RECs are undoubtedly considered a unique instrument for a just energy transition. There are two core features that define a REC:
  • RECs are the entities that should provide social, environmental and economic benefits to their members and respective local communities rather than solely for financial gain.
  • RECs should be effectively controlled by shareholders or members who are located in close proximity to the renewable energy projects owned and developed by REC’s legal entity (European Union, 2019, 2018).
Economic impacts of RECs can be various, ranging from effects on members and investors, on local and regional economies, and on electricity market and its stakeholders. Table 1 depicts a taxonomy of economic impacts of RECs. The pre-Directives literature (Bauwens et al., 2016; Brummer, 2018; Brummer et al., 2017; Candelise and Ruggieri, 2017; Heras-Saizarbitoria et al., 2018; Holstenkamp and Kahla, 2016; K.Huntala, 2016; Koltunov and Bisello, 2021; Kooij et al., 2018; Magnani and Osti, 2016; Magnusson and Palm, 2019; Tricarico, 2018; Vernay and Sebi, 2020; Wirth, 2014; Walker et al., 2010; Berka and Creamer, 2018; Moroni et al., 2019) mainly explored micro-scale socio-economic impacts on REC members with a specific focus on domains of management and sociology, while rarely examining the meso [^1] Pre-Directives, there are only few economic studies with those exploring RECs impact on the regional economies using input-output models (Lantz and Tegen, 2011; Okkonen and Lehtonen, 2016; Torgerson, 2006; Phimister and Roberts, 2012; Allan et al., 2011; Bere et al., 2015; Entwistle et al., 2014) and macro effects of RECs on the economy and markets.
In contrast to the multiple benefits at the organizational and local levels, REC deployment could lead to adverse effects from an economic perspective at macro level. For example, the extra injection of intermittent renewable energy could cause higher ramp/start events of expensive peaking plants, an increase in distribution grid charges due to RECs in the presence of volumetric tariffs could adversely affect households non-participating in RECs. Post-Directives, several economic empirical studies have appeared on the effect of RECs on the electricity system (Backe et al., 2022; Sarfarazi et al., 2020; Fuentes González et al., 2022; Di Silvestre et al., 2021; Boccard and Goetz, 2025). However, most post-Directive studies focus on the analysis of peer-to-peer trading, an innovative feature of REC that is yet rarely encountered in real-world applications [^2] More details on RECs classifications can be found in Rossetto et al. (2022), Kolesar (2022), Koltunov et al. (2025b) (Castellini et al., 2021; Chen and Gao, 2024; Dong and Li, 2024; Glachant and Rossetto, 2021; Hahnel et al., 2020; Hahnel and Fell, 2022; Nieto-Martin et al., 2019; Sousa et al., 2019).
Table 1: Taxonomy of REC economic impacts. Source: Koltunov and De Vidovich (2025).
Category Impact on members Impact on local Impact on market stakeholders
of impact and investors and regional economies and the electricity system
Scale Micro Meso Macro
Subject One EC Multiple ECs All RECs deployed
at a country level
Objects Individual members, Local and Generators, retailers,
investors (if not a member) regional economies DSOs, aggregators,
non-member consumers,
ESCOs, technology providers
The macro impact of REC deployment on electricity systems remains an underexplored domain (Koltunov and De Vidovich, 2025). Robinson and del Guayo (2022) propose two categories: systemic impact and stakeholder impact. The first category explores the impact of RECs on the electricity systems overall, which entails multiple spillovers on the operations of various actors. Examples of such systemic impacts include changes in distribution charges by Distribution System Operators (DSOs) (Berg et al., 2024), transmission charges by Transmission System Operators (TSOs), distribution and transmission system expansion, wholesale prices, changes in collected taxes and levies embedded in electricity tariffs, and market competitiveness. The second category examines the impact of RECs on specific stakeholder groups, such as non-members, retailers, generators, DSOs, aggregators, and technology providers. The present article belongs to the first category, as we focus on the merit-order effect and the impact of RECs on the Italian day-ahead (DA) power market equilibrium, which entails multiple spillovers on the operations of various stakeholders.
In terms of business model archetype, Italian RECs can be classified as a ’community collective generation’. [^3] In this type of REC, generation facilities must be connected to the same voltage substation and all members should live in its proximity, while individual members (households/SMEs/public entities/non-profit organisations) maintain their own retailers that take care of their residual demand. This type of REC is the one that is mostly adopted in Italy. This type of REC includes prosumers, consumers, and producers (i.e., the collective generation facility). During periods of excess of generation, RECs supply energy to the grid, thereby increasing renewable dispatch on the supply-side of the wholesale market. At the same time, these RECs aim for higher self-consumption, especially when the selling price is lower than the retail price, thereby decreasing the aggregated load on the demand-side of the market. For the scope of this paper, we state our research question as follows:
  • What is the impact of REC deployment on the Italian wholesale day-ahead market equilibrium?
Europe has the largest number of RECs, with almost 4,000 initiatives and 900,000 members (Koltunov et al., 2023). European RECs deserve particular attention because of the scale of the phenomenon and their formal recognition at the EU level by the RED-II (European Union, 2018) and IEM Directives (European Union, 2019), which triggered their further growth. Due to the lack of legislation and large heterogeneity of practices, RECs outside Europe remain a niche sector, thus restricting the potential for their deployment (Koltunov and De Vidovich, 2025). Within Europe, we selected Italian RECs as a case study for several reasons. First, they have shown substantial growth during last few years. [^4] In 2023, there were only 50 RECs aligned with new regulation in Italy (Koltunov et al., 2023). Currently, there are around 362555Data includes RECs in both operational and design phases as of January 2025 new RECs in Italy and the number continues to grow. The primary factors behind this notable deployment include the nuanced policy that allowed institutional actors to participate in RECs [^6] Historically, RECs were considered citizen-driven initiatives, which limited their deployment. and the premium tariff for plants of up to 1 MW capacity within a REC configuration (MASE, 2023). To date, the incentive for RECs is the only available subsidy for new photovoltaic installations in the country. Italy allocated €5.7 billion to support 5GW REC deployment: €3.5 billion for premium tariffs (financed via a levy on electricity consumption) and €2.2 billion from the National Recovery and Resilience Fund to cover up to 40% of CAPEX in small municipalities. These figures highlight the strategic importance of assessing the effectiveness and system-wide value of RECs. Second, the Italian wholesale day-ahead power market is well-suited for our research because of its transparency regarding hourly supply and demand bids placed by market operators. Third, Italian regulation is rather innovative and can be characterized as a “virtual scheme”, where REC members (prosumers and consumers) do not physically share energy in a microgrid. [^7] In contrast, CEER (2019) reports that “Some Member States, such as France and Austria, have developed a framework for collective self-consumption, where energy can be shared within a group of customers, without requiring direct involvement of a supplier.” Finland adopted a concept of energy sharing that is typically limited to apartment blocks or single housing associations. Similar to Finland, German “collective self-consumers” concept relates largely to occupants of the same building or small groups rather than a rule allowing multi-building or multi-substation virtual sharing across a distribution grid. Instead, they retain their existing retail contracts while sharing energy virtually, for which they are remunerated at the end of the year by the GSE [^8] GSE - in Italian, Gestore del Sistema Energetico - is the energy system manager, a public company that manages incentives in the energy sector on the state’s behalf. (Schiavo et al., 2022). This innovative regulation provides an interesting case for investigating its implications for the DA wholesale market. According to new rules (ARERA, 2022), REC members can benefit in four ways: energy self-consumption; a premium tariff for shared energy; valorization of avoided grid usage – calculated based on shared energy due to connection to the same primary substation; and energy sales (Blasuttigh et al., 2025). Figure 1 illustrates the benefits while Appendix 7.2 explains the formulas behind cashback components.
Notably, only the sales of injected energy (by prosumers and producers) and the self-consumed energy (by prosumers) affect the wholesale market, with the former impacting the supply-side and the latter the demand-side. The incentivized ‘shared energy’ is merely a virtual concept that does not directly trigger wholesale market changes. Therefore, we analyze only the operations of REC members that directly influence the market – namely, prosumers and producers – and disregard consumers. [^9] When joining a REC, consumers do not necessarily change their consumption profile in the ‘virtual scheme’ applied in Italy, therefore the wholesale market outcome does not amend because of them. Nonetheless, a REC manager/aggregator might advice consumers to change consumption in certain hours for the maximization of the common shared energy. Such advanced coordination requires smart infrastructure at the consumers’ premises and could potentially affect the wholesale market situation. However, to the authors’ best knowledge, no RECs in Italy have yet employed such level of coordination. Therefore, this aspect lies beyond the scope of the current study and remains for further investigation. From this perspective, our study pertains to the literature on the impact of prosumers on wholesale markets, although using Italian RECs as a specific case.
Figure 1: A net cash flow for REC members in Italy (excl. explicit capital subsidy).
We apply an innovative mixed-model methodology. In the first stage of analysis, we design a bottom-up engineering model that simulates the hourly behavior of various typical prosumers and producers within a REC. Then, we project the model’s results onto several scenarios, which are based on the policy target for 2027 and on an actual REC deployment status in 2024. The outcome of the first stage is then inserted into our synthetic model that simulates short-run schedule for the DA market on an hourly basis, from which we derive the effect of RECs on the merit order and aggregated demand.
Hence, the novelty of our work is twofold. First, we provide a theoretical contribution to the understanding of the seasonal and hourly impact of prosumers with different load profiles on the wholesale power market. We take Italian RECs as a specific case study. The heterogeneous membership within RECs becomes typical in many EU countries, especially after the implementation of the RED-II and IEM Directives. However, the existing literature on modeling prosumers’ behaviour mostly simulates the impacts of residential prosumers at a single geographical location, whereas our study simulates RECs as a collection of prosumers with very different consumption profiles (public, residential, small and medium enterprises, non-profit organizations) and standalone producers at various climatic locations, which closely reflects the country-scale REC deployment situation. Our contribution is tested empirically for the entire year 2024 in the DA Italian pool using a synthetic approach to simulate real-world competition among market agents. By undertaking such a detailed modeling exercise, our theoretical contribution may achieve higher external validity, especially for jurisdictions and markets with similar technical definition of a REC (e.g., Spain, Greece, Czech Republic), climate conditions and generation profiles (e.g., Spain, Greece).
Second, we provide a novel mixed-model approach to assess the country-scale impact of prosumers and producers aggregated into a REC on wholesale market scheduling. The first stage of our methodology begins with detailed data collection, which incorporates information on the composition of prosumers within RECs, the capacities of generating plants, the load profiles of different categories of prosumers, and the geographical distribution of RECs. Using these real-world parameters, we design the engineering model to obtain the assumed prosumers’ generation and load profiles, which we then project to the policy-targeted deployment level and other predicted deployment levels. In the second stage, the methodology is expanded by incorporating the previous outputs into an economic synthetic model. In essence, our methodological approach can be used by policymakers and stakeholders to determine the actual and predicted effects of the growing presence of heterogeneous prosumers, including, but not limited to, those arranged within RECs, on wholesale markets.
The remainder of the article is organized as follows. In Section 2, we explore the existing literature on systemic impacts of RECs and studies that utilized the ”synthetic approach” to modeling electricity markets. Section 3 explains the novel methodology employed in this paper, while Section 4 displays the major results. Section 5 incorporates the results into the academic debate. Section 6 concludes with final remarks and policy recommendations.
2 Literature review
This Section reviews the relevant literature on the topic, by exploring two main areas [^10] We also discuss the studies that analyze the effect of prosumers on electricity systems in Section 5. However, since they do not constitute the central role to our paper, unlike studies that focus on the effect of REC deployment on the system, we omit reviewing them in the main text although including them in the Appendix 7.4.:
  1. 1.
    Studies that analyse the effect of RECs on electricity systems;
  2. 2.
    Studies that employ a counterfactual approach for modelling the impact of specific power generation sources (e.g., RES) on the wholesale power market.
2.1 Impact of RECs on electricity systems
Most studies use a theoretical narrative-based approach to discuss the systemic impacts of RECs (Biggar and Hesamzadeh, 2022; Del Pizzo et al., 2022; Parag and Sovacool, 2016; Robinson and Del Guayo, 2022; Di Silvestre et al., 2021) and only few empirical economic studies exist (Backe et al., 2022; Sarfarazi et al., 2020; Fuentes González et al., 2022; Di Silvestre et al., 2021; Boccard and Goetz, 2025) due to the novelty and previously small scale of the phenomenon. Several studies indicate systemic benefits of REC deployment. Encouraging investment in embedded generation at a larger scale increases economic efficiency (Biggar and Hesamzadeh, 2022). Locating RECs in places with under-supply of electricity reduces transmission grid costs and losses (Robinson and Del Guayo, 2022). RECs reduce the need for an additional large-scale generation due to the merit-order effect, which in turn reduces the wholesale price and the need for additional transmission infrastructure (Robinson and Del Guayo, 2022; Backe et al., 2022). Robinson and Guayo (2022) argue that RECs can provide ancillary services if aggregated into virtual power plants, thereby replacing conventional sources which are more expensive to operate. When RECs are aggregated and utilize battery energy storage systems (BESS), they could reduce peak heat load and total heat demand in highly electrified systems (Backe et al., 2022). New ESCOs and aggregators could emerge as facilitators between RECs and the grid, therefore enhancing residential and commercial energy efficiency efforts (Parag and Sovacool, 2016). Finally, multiple social benefits (Biggar and Hesamzadeh, 2022; Robinson and Del Guayo, 2022) and improved competitive pressure on other market agents (Robinson and Del Guayo, 2022) are also important systemic benefits.
At the same time, other studies point to systemic detriments of REC deployment. Biggar and Hesamzadeh (2022) and Robinson and Guayo (2022) argue that a REC may arbitrate [^11] Arbitrage effect emerges when collective generators and consumers/prosumers have distinct selling and buying prices. When they unite into an energy community, generator now can supply energy directly to members of the REC while latter remunerate the former with a payment that is higher than conventional price for selling into the grid but lower than a conventional buying price for consumers. As a result, both parties win while the regulatory effect of the price has vanished. The buying price for consumers from local generators can be implicitly lowered compared to buying price from a central market if the distribution/transmission charge is distance-based. between retail and selling prices in cases when it acts as a united entity which purchases part of its electricity externally from a single retailer. As a consequence, the effect of non-uniform tariffs (time-of-use and locational tariffs) is eliminated, which could make the system operation costlier and potentially may lead to the further expansion of the network. Biggar and Hesamzadeh (2022) indicate that retail customers do not typically face the correct incentives to use energy according to market and grid situations. When customers unite into RECs it may further amplify the impact of such inefficient tariffs. The same study (2022) argues that the reduction in network exploitation exacerbates network operator’s ‘death spiral’ effect [^12] Death spiral effect describes a situation when distributed energy resources lead to prosumers paying less with volumetric distribution charges. Consequently, system operators (DSOs/TSOs) become under-paid for its previous infrastructure investment, which is usually higher than a decreased usage of a network due to self-consumption. System operators shift these costs instead onto other consumers, which in turn stimulate them to transform into prosumers as well, thereby shrinking the customer base iteratively.. Robinson and Guayo (2022) and Parag and Sovacool (2016) say that locating RECs in places with over-supply of electricity increases transmission grid costs and losses due to increased electricity injection where congestion bottlenecks are present. In addition, small capacity renewable generators are costlier for the system because they are more expensive by LCOE than large capacity renewable generators (Robinson and Del Guayo, 2022). Parag and Sovacool (2016) point that RECs might erode sensitive protections on privacy. As we can see, the literature points to multiple potential harms to the system and stakeholders [^13] Since our study aims to analyze systemic impacts, we skip the literature review of impacts on the individual stakeholders in the main text. Instead, we include the review of these impacts in Appendix C. from the deployment of RECs. However, these potential harms relate to a variety of REC types and not all are necessarily applicable to the ‘community collective generation’ type deployed in Italy [^14] Koltunov (2025a) can be consulted for an in-depth discussion on the impacts of individual REC business models on the electricity system and its stakeholders..
Fuentes Gonzalez et al. (2022) utilize a game theoretical approach with optimization modelling for a simplified three-nodal Chilean market. The study reveals that the nodal price decreases from 65.77 \$/MWh to 65.4 \$/MWh when RECs are deployed in a scenario with 9.24 MW of installed capacity with 150 MW peak nodal load. Moreover, the price drops further up to 64 \$/MWh if a REC’s installed capacity gets higher, though contingent upon the disposable income of REC members. The reduction in the nodal price suggests that the effect of RECs can also be positive on the non-member consumers. An increase in social welfare can be observed when RECs are involved in energy production (14,221 \$/hour with RECs versus 14,596 \$/hour without REC) due to the avoidance of high generation costs at nodes without REC, as well as the less frequent need for transmission expansion. This leads to a reduction in the nodal price that triggers an increase in the quantity demanded from 105.79 MWh to 115.03 MWh with an energy community. Quantities demanded in their static model are simulated for a single representative hour while prices are simply derived from linear demand functions. Consequently, this approach does not take into consideration a full interplay of technologies and variability over multiple time periods that are inherent in pay-as-clear zonal wholesale auctions of the real world.
Sarfarazi et al. (2020) also use a game-theoretical approach with optimization and dynamic programming to simulate the interaction of a stylized German REC with its retailer and a wholesale market. In the first scenario, a community storage is owned by the retailer who pursues the profit maximization objective. Here, real-time buying and feed-in selling tariffs do not decrease the quantities dispatched at the supply side of the wholesale market by RECs compared to static buying tariffs. In the latter case, community storage decreases the energy dispatched at the supply side of the wholesale market. In general, the retailer’s profit-maximization scenario leads to welfare gains both for the system and a REC. Conversely, if a retailer pursues a REC self-sufficiency maximization objective and not a profit maximization objective, REC imports less electricity from the public grid, therefore, less of grid charges are collected for the DSOs and less of levies to the general budget. Sarfarazi et al. (2020) focus primarily on the ”REC-retailer” interactions on a micro scale and do not investigate REC deployment on a macro scale, merit-order effect (MOE) and complexities of a DA wholesale market.
Boccard and Goetz (2025) find that the impact of RECs on the grid can be diverse. When a PV system is sized at 25% of the aggregated load of REC residential members the annual power exchanges with the grid drop from 70 MWh (no PV scenario) to 59 MWh (14% decrease) in Northern Spain. This is a significant finding, which indicates that a DSO could manage the deployment of RECs and generally prosumers without having to raise tariffs and compromising on grid reliability. When the PV capacity is increased to cover 50% of an aggregated load, the power exchanges with the grid decrease by 7%. Instead, in the scenario where a PV capacity covers 100% of the member load, power exchanges with a distribution grid raise up to 29%. The situation worsens in the northern irradiance region (Germany), whereas betters in the southern irradiance location (Portugal). In contrast, in the scenario with heterogeneous public prosumers (schools and hospitals) that install PV systems covering 50% and 100% of the load, the power exchanges reduce by 8% and 13%, respectively, compared to scenario without a PV. Instead, the 25% PV coverage of public loads rises power exchanges with the grid by 10%. The authors argue that the deployment of small-scale PV systems in residential RECs could decrease operational costs for the grid, while large-scale PVs force costly upgrades within residential buildings and across a distribution grid, and therefore better suit RECs composed of public prosumers.
Last, Backe et al. (2022) utilized two large optimization models to explore the effect of the REC deployment on the additional capacity needs in a 2060 carbon-neutral Europe. The authors find that the REC deployment reduces both electricity and heating costs and lessens the need for national capacity expansion by 50-60 GW in six EU countries. However, they do not quantify the MOE or REC interaction with the DA market, openly stating that this is beyond the scope of their paper.
2.2 Synthetic approach to simulating electricity markets
While optimization models only represent in detail the behavior of a firm, equilibrium and simulation-based models represent market behavior considering competition among all participants (Ventosa et al., 2005). Equilibrium models are based on the definition of the equilibrium which is mathematically expressed in the form of a system of algebraic or differential equations, which imposes limitations on the representation of competition between participants and are frequently too hard to solve (p. 5, 2005). Simulation models are an alternative to equilibrium models when the problem is too complex to be addressed by researchers within a formal equilibrium framework (p. 6, 2005). Therefore, simulation models provide a more flexible way to address the market problem than equilibrium models, which justify its usage for our research purpose. However, a limitation of the simulation models is that they are based on assumptions that are particular to each study.
The study by Sensfuß et al. (2008) is one of the first in which the “synthetic supply” approach was applied to analyze the impact of renewable electricity generation on the spot market equilibrium. This approach involves modeling a counterfactual scenario without generation from RES and comparing it with the actual scenario where generation from RES is present. The authors find that, between 2001 and 2006, the price reduction due to the merit-order effect from renewable generation in Germany was significant, reaching its peak in 2006 at approximately €5 billion. The paper concludes that the economic benefit of RES proliferation is greater than the cost of subsidies. The net profit for consumers was €1.9 billion in 2006 alone.
A prominent topic in the simulation-based literature is the measurement of market power. For example, Ciarreta and Espinosa (2010) apply a “synthetic supply” approach in the Spanish electricity market to show that actual market prices were about 21% higher than those in a counterfactual scenario without strategic bidding by large firms, especially in 2002 and 2005. Rossetto et al. (2019) conduct a similar exercise for the Italian market between 2015 and 2018, finding that consumer surplus losses attributable to the dominant operator grew over time and were most pronounced during peak-demand months.
Beltrami et al. (2021) examine the merit-order effect in the Italian market, demonstrating that RES displace conventional generators and thereby lower overall wholesale prices. Notably, by subtracting subsidies from the environmental and economic benefits, the study still finds a 44% net welfare increase in 2018 alone, which is similar to findings from the Sensfuß and Genoese (2006) for Germany. Espinosa and Pizarro-Irizar (2018) similarly employ a synthetic supply approach to estimate the social costs and benefits of Spain’s RES subsidies over the 2002–2017 period. They show that while the subsidies initially provided net social benefits, cuts to these subsidies eventually led to reduced RES deployment, which in turn diminished the merit-order effect and offset many of the gains. Turning to electricity storage, Beltrami (2024) investigates pumped hydro storage in Italy. The ‘synthetic’ simulation reveals that the CO2 saved during discharge exceeds the CO2 generated during charging.
Synthetic approach has also been applied to hypothetical market scenarios. Ciarreta et al. (2024) explore a prospective Moroccan electricity market, comparing two counterfactual market designs to a status quo scenario without a liberalized market. They find price reductions of 48% and 43%, largely driven by more efficient dispatch and interconnection usage with Spain, and emphasize the importance of block bids in lowering final prices.
3 Methods and data
The first stage of our methodological approach begins with mapping all Italian RECs to derive input parameters for subsequent engineering modeling. The model’s output is used to project REC deployment for a year 2027 according to three scenarios, which provides the input variables for the second stage. In the second stage, we first simulate a market equilibrium with RECs in 2027. Then, we construct a counterfactual scenario of market equilibrium without RECs. As the most recent annual data on day-ahead market bids is available only for 2024, and the policy target is set for 2027, we assume that the wholesale market conditions in 2027 are identical to those of 2024 for simplification purposes. This assumption does not compromise our modeling objective, which is to introduce a new methodology for estimating the impact of REC deployment on wholesale market equilibrium. Our aim is not to forecast the actual impact of REC deployment in the policy-target year. Nevertheless, policy-relevant insights can be drawn from the modeled scenarios. In the first stage of methodology, we used a combination of software applications: MS Excel (mapping), Matlab (engineering modeling), R (projection). In the second stage, R was utilized as the main modeling software.
3.1 First stage: Mapping, Engineering modeling, Scenarios and Projections.
3.1.1 Comprehensive Mapping of Italian RECs
To accurately model the actual and potential impact of RECs, we began by collecting data on all operational and planned RECs in Italy. [^15] The complete data on the mapping of Italian RECs are available in Supplementary materials. Due to the absence of a comprehensive dataset from a single source, we compiled our database from multiple sources, specifically:
  • the data portal of GSE [^16] The state-owned renewable energy agency,
  • the annual reports of Legambiente [^17] A national environmental non-governmental organisation, i.e. NGO.,
  • the data portal NeXt [^18] NeXt — the civil society network comprising the majority of Italian third-sector and public bodies working in the REC field ESG,
  • the data portal Sinergie Condivise [^19] Developed by a banking foundation ‘Fondazione Compagnia di San Paolo‘ in collaboration with regional and municipal governments and universities in the Piedmont Region,
  • publicly available business plans of individual RECs,
  • academic publications,
  • websites of REC developers,
  • websites of news agencies, regional/municipal authorities, and individual initiatives.
We refer to Italian electricity market zones using acronyms: North – NORD, Central North – CNORD, Central South – CSUD, Calabria – CALA, South – SUD, Sicily – SICI, Sardinia – SARD. The prosumer’s categories (public, residential, SMEs, NPOs) were determined based on subjects that are entitled to become a member of REC by Italian law (ARERA, 2022), namely public entities, private citizens, SMEs, and NPOs. Our final database contains 34 variables for 362 RECs in Italy, of which 184 are in the operational phase and 178 are in the design phase as of January 2025. However, for the purposes of this study, we used only 20 variables. Zhu et al. (2025) also built the database of new Italian RECs for the purpose of their study. In addition to the diverse scope of variables collected (due to different research objectives), another major difference lies in the development status of the RECs. Our database includes both operational RECs and those that were in the project design stage of development. Inclusion of RECs in the project design stage allowed us to have a more holistic perspective on future trends, which is discussed in further detail in Section 3.1.3. In turn, the database collected by GSE (2025b) though is complete but contains a very small set of variables  [^20] The recent map published by RSE (2025) is primarily based on the GSE data, although contain more variables than the original database, but still does not include critical variables (i.e. related to categories of members) used for our research.. However, neither of the other databases contain information about a category of members nor about the shares of installed PV capacity by prosumer categories. In contrast, our database contains these variables because it was collected specifically for the purpose of this study. A detailed comparison of three databases can be found in Table 3.1.1.
Comparison between the databases Variablea Our database Zhu et al. (2025) GSE (2025b) Total number of RECs, from which: 362 212 344 Operational RECs 184 212 344 Design-phase RECs 178 0 0 Location 362 212 344 Population (of municipality) 85 Climatic zone 85 Market zone 362 Generator’s capacity 300 212 344 Type of technology installed/planned 323 85 Number of membersb 283 344 Category of members 219 Shares of installed PV capacity by category of prosumers 271 Technical indicatorsc 147 BESS availability 13 7 EVs and/or e-charging availability 38 3 Self-consumption level 82 Economic indicatorsd 105 Number of prosumer buildings by type 158 55 Number of consumer buildings not counted 41 Investment source 252 CAPEX costs 106 Legal form 180 344 REC builders and promoters 204 DSO operating a primary cabin 344 Last update January 2025 February 2025 October 2025 Data sources GSE, Legambiente, RSE, Sinergie, Condivise, NeXt ESG, business plans, academic publications, websites of builders, other websites GSE, Legambiente, RSE, academic publications own
  • a
    Some reported variables represent a set of variables in actual database. For example, ”Shares of installed PV capacity by category of prosumers” contain five variables in the database, etc.
  • b
    Members include technical users (prosumers, consumers, producers), and non-user members.
  • c
    Technical indicators collected from business plans include energy generated, injected, and shared as well as self-consumption level.
  • d
    Economic indicators collected from business plans include revenues from energy sales and energy sharing, savings from energy self-consumed.
From our database, we derived input parameters for the subsequent methodological stages. The first input parameter can be calculated via the following equation:
$P_{p,z}^{PV,avg,one}=\frac{P_{p,z}^{PV,avg,total}}{N_{p}^{r}}$ (1)
where
  • $P_{p,z}^{PV,avg,one}$ is the average PV capacity per one prosumer/producer of category $p$ in electricity market zone $z$.
  • $P_{p,z}^{PV,avg,total}$ is the total average PV capacity of all prosumers/producers of category $p$ within an individual REC in market zone $z$.
  • $N_{p}^{r}$ is the average number of rooftop or standalone installations for prosumers/producers of category $p$ across the entire country [^21] Due to the small sample sizes for the number of rooftop/standalone installations across individual zones in our database, we decided to estimate the average number of rooftop/standalone installations across the entire country, that is, $N_{p}^{r}$. However, if the sample size allows, the methodology should instead employ the averages across individual zones, that is, $N_{p,z}^{r}$ .
  • $p$ denotes a category of prosumer/producer, where $\mathcal{P}\in\{\text{Public},\text{Residential},\text{SME},\text{NPO},\\ \text{Standalone producing installation}\}$.
  • $z$ denotes an electricity market zone, where $Z\in\{\text{NORD},\text{CNORD},\text{CSUD},\text{CALA},\\ \text{SUD},\text{SICI},\text{SARD}\}$.
  • oneone denotes one prosumer/producer.
In turn, the array of $P_{p,z}^{PV,avg,total}$ values was derived via the following equation:
$P_{p,z}^{PV,avg,total}=\frac{1}{n_{p,z}}\sum_{n=1}^{n_{p,z}}P_{p,z}^{PV,ind}\left(\frac{Sh_{p,z}^{ind}}{100}\right)$ (2)
where
  • $P_{p,z}^{PV,ind}$ is the PV capacity of all prosumers/producers of category $p$ within an individual REC located in market zone $z$.
  • indind refers to an individual REC.
  • $Sh_{p,z}^{ind}$ is the capacity share of prosumers/producers of category $p$ within an individual REC in zone $z$.
  • $n_{p,z}$ is the number of RECs that include at least one prosumer/producer of category $p$ in market zone $z$.
Similarly, the self-consumption level of 49.1% was identified as the average across all RECs. Consequently, our engineering model is based on a self-consumption range of 45%–50%–55%, where the central value corresponds to the real-world situation, the lower bound reflects a slightly more pessimistic scenario, and the upper bound represents a more optimistic scenario that reflects the potential deployment of BESS. However, only 3.6% of the RECs currently report having plans to install BESS. Although BESS would enable both a higher self-consumption rate and increased remuneration for shared energy, long payback periods inhibit their adoption.
Another input parameter is the ’most common building type’, which is essential to select realistic prosumer load profiles for the engineering model. For example, public prosumers may be represented by the load profile of a school or a sports facility, while SME prosumers may be modelled using the load profiles of commercial buildings or hotels. Clearly, different load profiles yield different results. Therefore, identifying the most common building type is critical for producing outputs that closely reflect reality without introducing excessive complexity by modelling all possible building types. This parameter was derived from our database through manual counting of building types for different categories of prosumers.
3.1.2 Bottom-up Engineering Modelling of RECs
Figure 2: Graphical representation of the prosumer energy modeling framework used to simulate hourly energy flows for different prosumer/producer categories. The flowchart illustrates the input data (PV production and load profiles), the modeling process for each prosumer/producer category (residential, school, commercial, office, standalone), and the generation of annual hourly datasets for energy production, self-consumption, and grid injection across multiple market zones and self-consumption scenarios. These outputs are then used as inputs for the scenario-based projection and economic simulations in the day-ahead electricity market.
Modeling RECs at a fine-grained level allows a detailed assessment of how different categories of users interact with their renewable energy systems over time. This insight is crucial for calculating quantities such as self-consumed energy and injected energy into the grid, which are useful for determining the change in energy volumes and thus for estimating economic market variables. An overview of the proposed bottom-up engineering methodology is shown in Fig. 2, which is explained in the following paragraphs.
To develop our behavioral energy model of prosumers, we designed a modeling scheme that represents five prosumer/producer categories, namely residential, schools, commercial, office, and standalone PV systems. Except for the standalone PV systems (which act purely as generators without any demand), all user categories are equipped with both PV production and electric load, connected to the national power grid.
From real-world data collected across operating RECs in Italy, as described in the previous Section 3.1.1, we estimated the average PV nominal capacity installed for each prosumer category within different RECs. These values were then used to scale hourly-based annual PV production profiles, obtained from location-specific PV yield profiles for seven representative cities covering the main Italian market zones: Milan for NORD, Florence for CNORD, Rome for CSUD, Brindisi for SUD, Catanzaro for CALA, Palermo for SICI, and Cagliari for SARD.
Consumption patterns for each prosumer were modeled using nPro (Wirtz, 2023), a load profiling tool capable of generating hourly demand curves based on user-specific behavior. To explore different energy scenarios, the load profiles were scaled to achieve three predefined self-consumption levels: 45%, 50%, and 55%. This approach resulted in the generation of annual hourly-based time series (8,760 values) for each prosumer/producer category and each scenario, capturing PV generation, electricity consumption, self-consumed energy and energy injected into the grid. With seven market zones and three levels of self-consumption considered, the final output of the engineering model consists of twenty-one structured datasets. Each dataset includes the energy behavior of the four prosumer categories and the standalone PV systems, whose entire production is fed into the grid. These profiles are then used as input for the subsequent projection step, which is described in Section 3.1.3.
Photovoltaic Generation Profiles
To characterize the hourly production of PV systems across Italy, we made use of the PVGIS platform developed by the European Commission’s Joint Research Centre (PVGIS, 2022). For each of the seven reference city, we extracted location-specific PV yield data. Rather than relying on generic assumptions, we configured each PV system in PVGIS under optimal operating conditions. This meant that, for each site, the tilt angle of the PV modules was set to the value maximizing annual output, the azimuth orientation was chosen to face due south, which typically yields the best yearly performance in Italy and the PV technology considered was crystalline silicon, the most widespread module type. Also, system losses, such as reflecting factors, inverter efficiency, temperature effects, cable losses, dust, and shading, were included using an average value of 15% (Ogliari et al., 2023).
For each location, the tool provides an hourly-based annual time series representing the energy yield per installed kilowatt-peak. These values were then scaled by the average installed PV capacity for each prosumer/producer category $P_{p,z}^{PV,avg,one}$, based on data from existing RECs, as in (3):
$E^{PV,one}_{p,z}(t)=Y^{PV,one}_{p,z}(t)\cdot P_{p,z}^{PV,avg,one}$ (3)
where
  • $E^{PV,one}_{p,z}(t)$ [kWh] is the hourly-based annual profile of energy produced by the PV system of category p in electricity market zone z for one prosumer/producer.
  • $Y^{PV,one}_{p,z}(t)$ [kWh/kWp] is the hourly-based annual yield profile of the reference city from PVGIS of category p in electricity market zone z for one prosumer/producer.
  • t is the time index $\in\{1,2,...,8760\}.$
By using PVGIS in this way, we were able to obtain consistent, reproducible photovoltaic production profiles that reflect regional climatic differences and common technical configurations, without introducing excessive complexity into the modeling process. It’s worth noting that, while real-world data from monitored PV systems might offer greater accuracy in principle, such datasets are often fragmented, inconsistent, or not openly available across all regions. Simulated data from PVGIS, on the other hand, ensures full spatial coverage and comparability, while still being grounded in satellite-based irradiance data and validated performance models.
Load Consumption Profiles
To represent electricity consumption behavior for each prosumer category within the REC framework, we generated hourly demand profiles using nPro (Wirtz, 2023), a profiling tool that provides synthetic yet behaviorally-informed load curves. The tool produces time series based on statistical models of daily and seasonal usage patterns for different consumer categories (residential, schools, commercial, office, and many others) under typical operational conditions  [^22] All prosumers are assumed to own electrical loads related to heating, cooling, and general electricity demand, except for residential prosumers, for which only cooling has been considered in order to model summer air-conditioning. For all loads, the default values provided by nPro have been retained, meaning that no additional calibration of consumption profiles was performed beyond the standard dataset assumptions. . However, while these base profiles are useful in capturing the temporal distribution of consumption during the year, they do not by default reflect a specific relationship with local PV generation. In particular, no predefined level of self-consumption (i.e., the portion of PV energy that is immediately consumed by the user) can be assumed unless demand and generation are explicitly aligned.
To introduce variability in this key parameter and explore its influence on community-level energy flows, we implemented a simple but effective adjustment strategy. The idea was to scale the demand profiles vertically to increase their overall magnitude, without altering their temporal pattern. This approach changes the extent to which demand and generation tend to coincide and allows us to control the self-consumption ratio, a value that is typically chosen when designing PV systems, in straightforward way.
Specifically, for each user category p in electricity market zone z, the original load profile generated by nPro was multiplied by a constant factor which was iteratively tuned until the ratio of self-consumed energy to total PV production matched a given target. We considered three such targets (45%, 50% and 55%) chosen to represent a plausible range of self-consumed energy. In this setting, energy not consumed at the time of generation is immediately injected into the grid, as no energy storage system is supposed to be available.
However, to account for the likely evolution of REC configurations in the coming years, we also explored additional scenario featuring higher self-consumption levels (mixed scenario in Table 2). This extended case is intended to emulate the effect of widespread adoption of energy storage systems, such as residential or commercial batteries. By shifting consumption toward daylight hours or enabling the deferred use of solar generation, energy storage reduces the amount of surplus energy injected into the grid, thus increasing the local use of renewable electricity. Although modelling of storage systems is not explicitly included in this framework, this scenario of increased self-consumption provide an indication of their aggregate effect on energy flows. This allows an exploratory assessment of how the progressive penetration of storage technologies may alter the energy balance and the economic impact of RECs on the market.
Calculation of Self-Consumed and Exported Energy
Once both the hourly PV generation and the electricity demand profiles were established for each prosumer/producer category and scenario, the next step involved calculating the two profiles that define the interaction between local generation and the grid: self-consumed energy and surplus energy injected into the main grid. These quantities were derived directly from the hourly time series of PV production $E^{PV,one}_{p,z}(t)$ and electrical load $E^{L,one}_{p,z}(t)$ of category p in electricity market zone z computed as explained in the previous sub-section. In particular, the self-consumed energy $E^{self,one}_{p,z}(t)$ of category p in electricity market zone z for one prosumer/producer at each hour is computed as in Eq.(4):
$E^{self,one}_{p,z}(t)=min(E^{PV,one}_{p,z}(t),E^{L,one}_{p,z}(t))$ (4)
This corresponds to the portion of the PV production that is immediately used to meet on-site demand. When the load exceeds the available PV power, all PV generation is consumed locally. Conversely, if generation exceeds demand, only part of it is self-consumed. In this case, the surplus energy exported to the grid $E^{exp,one}_{p,z}(t)$ of category p in electricity market zone z is calculated as in (5):
$E^{exp,one}_{p,z}(t)=max(E^{PV,one}_{p,z}(t)-E^{L,one}_{p,z}(t),0)$ (5)
In other words, the exported energy is the excess generation that is not used locally and is therefore injected into the public distribution network. These expressions were applied element-wise over the full annual time series, resulting in two additional vectors of hourly values for each prosumer/producer and scenario.
In the end, for each prosumer category, the combined simulation produces four time series per user: hourly PV production, electricity consumption, self-consumed energy and energy fed into the grid. Instead, for standalone PV producers, the energy injected into the grid is equal to the PV energy produced, hour-by-hour.
Figure 3: Hourly energy profiles for Milan (NORD) over a full year for four prosumer user categories. Each subplot shows PV generation (yellow line), electricity consumption (blue line), and energy injected into the grid (green line).
Figure 3 shows an example of the time profiles explained in the previous paragraphs for Milan (NORD). In particular, the annual hourly energy profiles for four different categories of prosumer users are presented. In each subplot, three key variables are shown over the full time horizon of one year. The yellow curve represents the electricity generated by the PV system, while the blue line indicates the hourly electricity consumption associated with the user. The green line, plotted below the horizontal axis, corresponds to the surplus energy that is not self-consumed and is instead injected into the grid. To make the graphs easier to read, the hourly profile of self-consumed energy is intentionally not shown.
The energy behavior varies significantly across prosumer categories. Residential prosumers exhibit relatively low and stable generation throughout the year, with a load profile that allows only limited self-consumption, especially during daytime hours. School buildings show a highly intermittent demand, with pronounced reductions in summer months and strong peaks during the winter, reflecting the academic calendar. Commercial prosumers show a more consistent and uniform demand profile, enabling a better alignment between PV production and consumption. Also, office buildings exhibit similar patterns, although with lower overall demand and more visible weekday-weekend variability. In all cases, the amount of feed-in is visibly related to the time mismatch between photovoltaic production and load demand.
Figures 4 and 5 display the same hourly profiles albeit over an average day in January and April. We see that in January (Figure 4) the greatest self-consumption (intersection between yellow and blue bars) is exhibited by a school. Moreover, its load (blue bars) surpasses its production during midday off-peak hours (10 a.m.–15 p.m.), thereby potentially offsetting an extra generation from REC’s residential and commercial prosumers. The load profile of an office is very similar to the school’s profile, though on a smaller scale (vertical axes). In April, we see that the absence of heating and cooling needs entails a more moderate demand for energy. Consequently, a lot of excessive energy goes into the grid. In both graphs, we observe that residential prosumers exhibit much smaller energy flows and a different load profile than other prosumers. In addition to evening peak difference, a peak midday load of a school, commercial and office buildings occurs at 9 a.m.–12 a.m., whereas of a residential building at 13 p.m.–14 p.m., although more pronounced in January than in April.
Figure 4: Hourly energy profiles for Milan (NORD) over an average day in January for four prosumer user categories. Each subplot shows PV generation (yellow bars), electricity consumption (blue bars), and energy injected into the grid (green bars).
Figure 5: Hourly energy profiles for Milan (NORD) over an average day in April for four prosumer user categories. Each subplot shows PV generation (yellow bars), electricity consumption (blue bars), and energy injected into the grid (green bars).
3.1.3 Scenarios and Projection of REC Deployment
The core objective of the simulation framework is to evaluate the market impact of RECs under a range of realistic deployment pathways. The calibration of the scenarios is primarily based on the fulfillment - or lack thereof - of Italy’s policy target of 5 GW of installed REC-linked RES capacity by year 2027 as well as on the current REC deployment levels (367.514 MW of PV capacity in operational and design status; 29.646 MW in operational status). Three main narrative trajectories are hereby defined:
  • Policy Scenario: this scenario assumes the full achievement of the 5 GW target by 2027, reflecting an optimistic rollout of RECs in terms of both regulatory support and investment mobilization.
  • Half-way (HW) Scenario: we assume the steady proliferation of the 2024 deployed capacity, with 0.368 GW in operational and design status, reaching 1.47 GW by 2027.
  • Business-as-usual (BU) Scenario: this scenario assumes no significant acceleration in REC deployment beyond the current trend, with 29.646 MW in operational status in 2024. For the 2027 projection, we again assume the steady proliferation of the 2024 operational capacity, reaching 0.119 GW.
Table 2: List of simulated scenarios
No. Scenario name Scenario code Year % of sc assumed GW of REC installed capacity Policy target achieving
1 Policy sc45.2027 2027 45% 5 Yes
2 Policy sc50.2027 2027 50% 5 Yes
3 Policy sc55.2027 2027 55% 5 Yes
4 Half-way sc45.HW.2027 2027 45% 1.47 No
5 Half-way sc50.HW.2027 2027 50% 1.47 No
6 Half-way sc55.HW.2027 2027 55% 1.47 No
7 Business-as-usual sc45.BU.2027 2027 45% 0.119 No
8 Business-as-usual sc50.BU.2027 2027 50% 0.119 No
9 Business-as-usual sc55.BU.2027 2027 55% 0.119 No
10 Mixed scenario sc_mix1.2027 2027 Public: 50% Residential: 45% SME: 55% NPO: 50% Standalone: no sc No
Within each of these trajectories, scenarios are further differentiated based on assumed self-consumption rates, defined as the share of generated renewable energy consumed by prosumers within a REC rather than injected into the grid. This parameter is a critical driver in the modelling framework, as it reflects the efficiency and “virtuosity” of RECs in optimizing on-site energy use - thus affecting both demand-side and supply-side market dynamics. Additionally, a “mixed scenario” is included to reflect more granular assumptions based on prosumer-category segmentation, as identified through our empirical mapping of REC profiles. These scenario applies differentiated self-consumption rates by prosumer category, thereby enhancing the realism and policy relevance of the simulated outcomes. A summary of all scenarios is provided in Table 2.
The scenario-based projection inputs were drawn from (i) the engineering model, which provided two key variables for one typical prosumer/producer within a REC—annual energy self-consumed and energy injected into the grid—and (ii) projection parameters derived from our database. The first projection parameter, “Zonal Shares of RECs,” represents the percentage distribution of RECs across electricity market zones:
$Sh_{z}^{total}$ $=\frac{100}{n}\sum_{n_{z}=1}^{n_{z}}n_{ind}$ (6)
where
  • $Sh_{z}^{total}$ denotes the zonal share of RECs,
  • $n_{ind}$ is an individual renewable energy community,
  • $n_{z}$ is the total number of RECs in each zone $z$, where $z\in Z$,
  • $n$ is the total number of RECs in Italy, $n=362$.
The “average capacity share of all REC prosumers/producers” is the second projection parameter. It is calculated using the following equation:
$Sh_{p,z}^{avg}$ $=\frac{1}{n_{p,z}}\sum_{n_{p,z}=1}^{n_{p,z}}Sh_{p,z}^{ind}$ (7)
where
  • $Sh_{p,z}^{avg}$ is the average capacity share of all REC prosumers/producers of category $p$ in market zone $z$.
  • $Sh_{p,z}^{ind}$ is the capacity share of prosumers/producers of category $p$ within an individual REC in zone $z$.
  • $n_{p,z}$ is the number of RECs where a prosumer category $p$ is present in zone $z$
The third parameter, ”total REC deployed PV capacity per prosumer category $p$ and zone zz” was calculated for each scenario using the equation:
$P_{p,z}^{PV,total}$ $=\frac{P_{scen}^{PV,total}\times Sh_{z}^{total}\times Sh_{p,z}^{avg}}{100}$ (8)
where
  • $P_{p,z}^{PV,total}$ is the total REC deployed PV capacity per prosumer category $p$ in market zone $z$,
  • $P_{scen}^{PV,total}$ is the scenario-based total REC deployed PV capacity.
The fourth parameter, the ”number of PV plants per prosumer category $p$ in market zone zz” can be calculated using:
$Plant_{p,z}$ $=\frac{P_{p,z}^{PV,total}}{P_{p,z}^{PV,avg,one}}$ (9)
where
  • $Plant_{p,z}$} is the number of PV plants per prosumer category $p$ in market zone $z$.
  • $P_{p,z}^{PV,avg,one}$} is the average PV capacity installed by one REC per prosumer category $p$ in market zone $z$.
By taking the outputs of the engineering model - $E_{p,z}^{\mathrm{exp,one}}(t)$ and $E_{p,z}^{\mathrm{self,one}}(t)$ - and knowing $Plant_{p,z}$, we can derive the projected energy injected into the grid and the projected energy self-consumed by different categories of prosumers/producers:
$E_{p,z}^{exp}(t)$ $=E_{p,z}^{\mathrm{exp,one}}(t)\times Plant_{p,z}$ (10)
$E_{p,z}^{self}(t)$ $=E_{p,z}^{\mathrm{self,one}}(t)\times Plant_{p,z}$ (11)
where
  • $E_{p,z}^{exp}(t)$ is the projected energy injected into the grid during hour $t$ by all photovoltaic plants of prosumer category $p$ in market zone $z$,
  • $E_{p,z}^{\mathrm{exp,one}}(t)$ is the energy injected into the grid during hour $t$ by one photovoltaic plant of prosumer category $p$ in market zone $z$,
  • $E_{p,z}^{self}(t)$ is the projected energy self-consumed during hour $t$ by prosumer category $p$ in market zone $z$,
  • $E_{p,z}^{\mathrm{self,one}}(t)$ is the energy self-consumed during hour $t$ by one prosumer of category $p$ in market zone $z$,
  • $t$ is a specific hour in a year, $t\in T$ and $T=\{1,2,\ldots,8760\}$.
Finally, we derive two arrays of projected variables: (i) energy injected into the grid by RECs, and (ii) energy self-consumed by RECs, using the following equations:
$E_{z}^{exp}(t)$ $=\sum_{p\in\mathcal{P}}E_{p,z}^{exp}(t)$ (12)
$E_{z}^{self}(t)$ $=\sum_{p\in\mathcal{P}}E_{p,z}^{self}(t)$ (13)
where
  • $E_{z}^{exp}(t)$ is the projected energy injected into the grid by all RECs during hour $t$ in market zone $z$,
  • $E_{z}^{self}(t)$ is the projected energy self-consumed by all RECs during hour $t$ in market zone $z$.
3.2 Second stage: Economic modeling
To assess the short-run economic impact of RECs on the Italian wholesale power market, we apply the empirical hourly counterfactual simulation proposed by Beltrami et al. (2021), further extended by Beltrami (2024). This methodology builds on publicly available data from Gestore del Mercato Elettrico – GME – namely, the “Offerte Pubbliche” – which report all price–quantity bids submitted to the Italian day-ahead auction. These data allow us to reconstruct the hourly merit-order demand and supply curves for each electricity market zone.
For a given hour and zone, each supply bid is characterized by an offered quantity $q_{i}^{s}$ and an associated bid price $p_{i}^{s}$, while demand bids are defined by requested quantities $q_{j}^{d}$ at bid prices $p_{j}^{d}$. Following the merit-order principle, supply bids are ranked in ascending price order and aggregated into a stepwise supply curve:
$S_{actual}(Pr)=\sum_{i:p_{i}^{s}\leq Pr}q_{i}^{s}$ (14)
Analogously, demand bids are ordered in descending price order and aggregated into a stepwise demand curve:
$D_{actual}(Pr)=\sum_{j:p_{j}^{d}\geq Pr}q_{j}^{d}$ (15)
By construction, both curves are piecewise-constant functions, reflecting the discrete nature of bidding in day-ahead auctions. The market-clearing price $Pr_{actual}$ is defined as the minimum price such that:
$S_{actual}(Pr_{actual})\geq D_{actual}(Pr_{actual}),$ (16)
with the corresponding cleared quantity given by $Q_{actual}=D_{actual}(Pr_{actual})$.
From now onwards, to maintain the focus on the synthetic counterfactual approach, we employ a simplified notation consistent with Subsection 3.1 In particular, $\Delta Q_{\text{REC,d}}$ corresponds to $E_{z}^{self}(t)$, while $\Delta Q_{\text{REC,s}}$ corresponds to $E_{z}^{exp}(t)$ in Eq. (12) and (13). [^23] Time and zonal indices are omitted for expositional clarity, while the strategy applies to all zones and settlement periods. Crucially, the counterfactual simulation aims to reproduce the hypothetical configuration of the day-ahead market in the absence of REC-driven generation and self-consumption. In this framework, the observed demand and supply curves, $D_{actual}(Pr)$ and $S_{actual}(Pr)$, constitute the empirical baseline from which synthetic curves are constructed to evaluate the equilibrium effects induced by RECs.
As stated above, RECs typically self-consume a share of their electricity production. This self-consumed electricity is not visible in the observed market demand. In a counterfactual scenario where RECs are absent, this hidden demand would need to be satisfied by the wholesale market. Let then $\Delta Q_{\text{REC,d}}$ denote the total quantity of REC self-consumption. To account for this, the demand curve is adjusted by shifting it horizontally to the right:
$D_{synt}(Pr)=D(Pr)+\Delta Q_{\text{REC,d}}$ (17)
On the supply side, RECs inject electricity into the market through surplus generation [^24] The GSE is responsible for informing the market operator, GME, about distributed generation volumes. The REC-driven quantities are then placed on the supply curve at zero price, along with the other distributed renewable generation.. Removing RECs implies that this contribution is also removed from the market supply. Let $\Delta Q_{\text{REC,s}}$ represent the total quantity of REC-generated electricity that would have been offered to the market. The adjusted supply curve is thus defined as:
$S_{synt}(Pr)=S(Pr)-\Delta Q_{\text{REC,s}}$ (18)
This corresponds to a horizontal shift of the supply curve to the left.
The counterfactual equilibrium price and quantity in the absence of RECs are denoted by $Pr_{synt}$ and $Q_{synt}$, respectively. These are determined by the intersection of the adjusted demand and supply curves:
$D_{synt}(Pr_{synt})=S_{synt}(Pr_{synt})$ (19)
Substituting the shifted functions, the equilibrium condition becomes:
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}=S(Pr_{synt})-\Delta Q_{\text{REC,s}}$ (20)
This can be rearranged as:
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}+\Delta Q_{\text{REC,s}}=S(Pr_{synt})$ (21)
This framework captures the dual effect of RECs on market equilibrium (see Figure 6). The removal of self-consumption by RECs increases observed demand, while the removal of injection by RECs reduces available supply, thus producing a rightward shift in the demand curve and a leftward shift in the supply curve. As a result, the counterfactual equilibrium price $Pr_{synt}$ is expected to be higher than the baseline price $Pr_{actual}$. The change in equilibrium quantity, $Q_{synt}-Q_{actual}$, depends on the relative elasticities of supply and demand. This formalization enables a quantitative evaluation of the role of RECs in lowering prices and reducing market dependency in wholesale electricity auctions.
Specifically, we modify the observed merit-order curves as follows:
  • Supply curve shift. In hours where RECs inject renewable electricity into the grid, we assume that this volume would not be available under the counterfactual scenario. Accordingly, the supply curve is shifted leftward to reflect the reduction in total market supply, primarily from RES. This leads to a counterfactual configuration in which electricity prices would be higher, ceteris paribus.
  • Demand curve shift. In hours where RECs self-consume a portion of their generation, the equivalent electricity demand is effectively removed from the market. In the counterfactual scenario, where self-consumption does not occur, we expand the demand curve rightward (upward shift in the merit-order framework), capturing the higher residual market demand that would otherwise materialize.
Figure 6: Graphical representation of the market equilibrium taking into account the effect of RECs. The curves $D_{\text{actual}}$ and $S_{\text{actual}}$ represent the actual demand and supply with RECs, leading to a market equilibrium at price $Pr_{\text{actual}}$ and quantity $Q_{\text{actual}}$. The dashed curves $D_{\text{synt}}$ and $S_{\text{synt}}$ refer to synthetic demand and supply, resulting in an alternative equilibrium at price $Pr_{\text{synt}}$ and quantity $Q_{\text{synt}}$.
This dual (contemporaneous) adjustment is performed hourly and for each market zone, generating synthetic demand and supply curves representing a no-REC baseline condition. The resulting counterfactual quantity outcomes are then compared with the observed market outcomes to estimate the impact of REC operations. To quantify the magnitude of REC effects on market outcomes, we define the hourly percentage impact on equilibrium quantities as the relative difference between the observed equilibrium quantity in the presence of RECs and the counterfactual equilibrium quantity in their absence. Formally, for each hour and market zone, the indicator is computed as:
$\text{Impact}_{Q,h}=\frac{Q_{actual,h}-Q_{synt,h}}{Q_{synt,h}}\times 100,$ (22)
where $Q_{actual,h}$ denotes the cleared quantity observed in the day-ahead market with RECs, while $Q_{synt,h}$ represents the synthetic equilibrium quantity obtained under the no-REC counterfactual scenario.
A positive value of the computed impact in Eq. (22) indicates that REC deployment is associated with higher market-cleared volumes relative to the counterfactual baseline, whereas a negative value signals a reduction in traded quantities, typically driven by higher levels of REC self-consumption. This metric is computed on an hourly basis and then analyzed across months, seasons, and policy scenarios, as reported in Section 4.
We acknowledge that this simulation is based on several assumptions. First, the counterfactual is computed under a ceteris paribus condition, assuming all other market dynamics unchanged. [^25] This assumption is particularly relevant when comparing the effects of RECs on demand and supply across the different scenarios, and becomes evident mostly in interpreting the results from the Policy Scenario compared to BU and HW. In our setting, we implicitly assume that the installed capacity of other RES technologies (beyond RECs) remains fixed. This does not reflect the actual dynamics of the Italian power system, where RES capacity additions are currently progressing at sustained rates as well as does not reflect the demand growth. Nevertheless, this choice is consistent with our primary goal, i.e. to isolate the specific marginal contribution of REC deployment, without counfouding their individual effect with broader renewable expansion trends. Second, each market zone is treated as a “closed system”, abstracting from inter-zonal electricity flows. While this assumption may limit economic and policy implications, it remains defensible for two main reasons: (1) RES-generated power is typically prioritized in dispatch due to the merit-order principle, making its theoretical removal analytically legitimate; (2) the current scale of REC operations in Italy is still limited, rendering its impact negligible in terms of strategic bidding by large market players deeply affecting the market clearing dynamics. Overall, this synthetic control approach allows us to isolate and quantify the localized, short-term effects of REC deployment on market outcomes such as zonal traded quantities and clearing prices, thus offering a robust empirical basis for evaluating REC-driven welfare gains.
4 Results
4.1 Descriptive statistics of REC configurations in Italy
As we see in Table 3.1.1, the information on the market zone and the type of installed technology is the most complete (362 and 323 RECs respectively). Crucial data on the nominal capacity of power plants is available for 82.9% (300) of RECs. The estimation of many parameters required data on the share of installed PV capacity by category of prosumers, which is available for 74.9% (271) of RECs. The typology of prosumer buildings is available for 43.7% (158) of RECs. In turn, the self-consumption level—from which the range of 45–50–55 scenarios was assumed—is available for 22.7% (82) of RECs. Similarly, only 3.6% (13) of RECs plan to install BESS.
Figure 7 shows that most RECs are situated in the NORD market zone (63.8%) and the Central-South zone (19.1%). In contrast, RECs are scarcely present in the rest of Italy. This deployment pattern may be associated with the general distribution of economic activity across the country. Since the second half of the 20th century, Northern Italy and the regions surrounding Rome have exhibited high levels of industrial and entrepreneurial activity. Consequently, the greater availability of expertise and financial resources for REC establishment has supported their rapid proliferation in these two market zones.
Figure 7: Zonal shares of RECs, $Sh_{z}^{\text{total}}$.
Figure 8: Distribution of RECs across prosumer/producer categories
In Figure 8, [^26] Figure 6 contains information not for all 362 RECs from our database but for RECs with the available information. we observe that the most common prosumer/producer category is public prosumers—typically represented by municipal buildings. Standalone systems are reported for 61 RECs, while rooftop installations owned by SMEs and residential prosumers are reported for 42 and 36 RECs, respectively. In contrast, only 22 RECs report involving NPO prosumers. The higher number of RECs with public prosumers may be explained by the fact that municipalities are, by far, the most common promoters of RECs in Italy, for whom targeted capital and regional grants are also more accessible. Moreover, public buildings consume most of their load during daylight and off-peak hours (see Fig. 4, 5), making them ideal candidates for maximizing the state incentive on shared energy (see Appendix 7.2). Accordingly, standalone plants also allow for the maximization of profits due to their greater average installed capacity, especially when balanced with the daytime loads of public, SME, and NPO prosumers and consumers. Although residential prosumers have not yet actively participated in RECs with rooftop systems, many residential consumers [^27] REC consuming members are not the focus of this study. Therefore, even though data on REC consumers is available, we do not report it here. are members of RECs in Italy [^28] The elaborate discussion on equity concerns and private citizen participation in Italian RECs can be found in Koltunov (2025b)..
Figure 9: Average capacity shares of all categories of prosumers/producers across market zones, $Sh^{a}_{p,z}$
Figure 9 shows the average capacity shares of all categories of prosumers/producers across market zones. Photovoltaic installations of all categories are present only in three zones: NORD, CSUD, and SUD. The CSUD and NORD market zones have the most proportionate prosumer/producer capacity distributions. This situation is possibly associated with the high number of RECs located in these zones (see Fig. 10). Moreover, the strong economic activity in these regions may contribute to a greater diversity of stakeholders who possess the financial, administrative, technical, legal, and social engagement resources and skills necessary to participate in RECs (Koltunov, 2025b; Musolino et al., 2023). In contrast, a more disproportionate distribution of capacity across prosumer/producer categories—where public bodies dominate—can be observed in the market zones of Southern Italy (SUD, SARD, SICI, CALA). Importantly, the absence of installed plants for certain member categories does not automatically exclude them from REC membership. All categories of members may also participate in RECs as consumers without owning rooftop systems.
Figure 10: Average PV capacity per one Prosumer/Producer, $P^{a.one}_{p,z}$.
Note: The number after the zone name indicates the sample size (REC observations).
Figure 11: Building/installation types by prosumer/producer categories.
Note 1: The numbers after building types indicate the number of individual buildings.
Note 2: “Others” in the Public category include police stations, autodromes, waste management companies, cemeteries, public utility facilities, etc. “Others” in the NPO category include social canteens, sports and educational centers, and social farms.
Figure 10 demonstrates the “Average PV capacity per one Prosumer/Producer“, $P^{a.one}_{p,z}$, which is one of the two real-world input parameters for the engineering model. The average capacity of a single photovoltaic plant varies across market zones. For public and SME prosumers, the average capacity in the NORD and Central zones (CNORD, CSUD) is significantly higher than in the Southern zones (CALA, SUD, SICI). [^29] Only 5 observations were used to estimate the average SME capacity in SARD. Therefore, 334.5 kW of average capacity installed by SMEs in SARD is based on a limited sample and should be interpreted with caution. Similarly, the very large capacity of public prosumers in CNORD zone, 285.5 kw, may be related to the small observational sample. Another notable observation is that NPOs have significantly smaller installed capacities on average compared to SMEs and public prosumers. Typically, non-profit organizations have more limited access to private financing than SMEs, while state capital subsidies are reserved exclusively for small municipalities (Koltunov, 2025b). Finally, zero values are mostly observed in the Southern zones, possibly due to the small number of REC observations.
In Figure 11, we observe the types of buildings—where rooftop installations have been constructed—estimated for prosumer categories. This information was used to identify the “most common building type.” For public prosumers, data on specific building types is unavailable for most RECs; consequently, the most common identified building type is a school. For SMEs, it is a commercial building, typically a supermarket or shopping mall. Hotels and industrial buildings are also relatively common in the SME category. The type of almost all residential buildings is unspecified (935 buildings), followed by private detached houses (14 buildings). Offices and churches are the most common building types for the NPO category. Finally, 65 standalone photovoltaic plants have been reported, while non-photovoltaic technologies have been used in a much smaller number of facilities.
We also analyzed the average number of buildings participating as prosumers in a single REC, $N^{r}_{p}$. On average, 4.4 public buildings with rooftop plants participate in a REC. Only 2.3 SME buildings and the same number of NPO buildings participate as prosumers in a REC. In contrast, approximately 17.7 residential houses participate with rooftop systems in a REC. However, as shown in Figure 10, residential prosumers have, on average, much smaller generating capacities. The average number of standalone producers per REC is 1.7.
4.2 Results from the synthetic approach
This subsection displays the core results of the paper, presenting the main outputs in weekly, monthly, and quarterly resolutions. The results are reported exclusively for the NORD and CSUD zones. [^30] Our simulation algorithm did not retrieve valid results for the remaining five zones, mainly due to data discontinuities and the low diffusion of RECs in such zones, and therefore we did not include them in the paper. Nonetheless, these results can be available upon request to the corresponding author. The analysis focuses on the comparative dynamics between observed market outcomes and their synthetic counterfactuals, highlighting the extent to which REC-driven injections and self-consumption might influence zonal market equilibria in terms of traded volumes and prices. To improve the readability of the plots that represent the percentage impact of RECs on equilibrium quantities in the DA market, we applied a smoothing procedure to all Figures from 16 to 25 using a centered simple moving average. Specifically, for each hour of the day, we replaced the original values with the average computed over a 7-hour symmetric window, using the rollmean function from the zoo package in R. The smoothing was applied selectively, only to positive values, while non-positive entries were left unchanged. This method reduces short-term fluctuations and highlights general trends across the different scenarios and seasons.
4.2.1 Results for main scenarios
[…]
In Figure 12, the notable difference between the magnitude of the equilibrium effects in winter (smaller spike density) and other seasons (greater spike density) can be observed. A combination of the low solar irradiance while high-self consumption in winter induces this trend. First, lower solar irradiation in winter lead to fewer quantities of energy to be injected into the grid by RECs. Second, self-consumption rate by RECs is higher during winter. Figure 13 illustrates daily self-consumption levels for all prosumer categories in different months. All categories self-consume on average more energy in January than in April, July or October. As a result, the RECs’ self-consumption effect on the demand side of the market is relatively bigger than the RECs’ injection effect on the supply side of the market in winter than in other seasons. In our engineering model, we added electrified heating and cooling to the load profiles of public (schools), SME (commercial), and NPO (offices) categories [^34] However, residential prosumers own just electrified cooling systems in our model and not heating systems, similar to the actual status quo in Italian households.. Therefore, in Figure 13, we observe that heating needs increase self-consumption during winter, mostly, for only three, albeit REC dominating (Figure 9), categories of prosumers.
Figure 12: Hourly percentage impact on equilibrium quantities from deployment of RECs in the North market zone.
Figure 13: Daily self-consumption rates for all prosumer categories in 12 months of the year.
[…]
By focusing on the NORD zone and delving into seasonal variations, Figure 15 displays the average profile of actual and counterfactual market quantities by hourly settlement period for January and April 2024. Under the HW scenario assumption, RECs deliver a net positive effect on actual market quantities only for central hours of the day in January (left panel). In contrast, during a typical spring month (April), the positive effect on market quantities is already visible since 6am, lasting until 7.30pm, in line with the fading sunlight (right panel). Thus, the scale of the impact of RECs is, as expected, stronger in April, because of the larger penetration of RES from RECs and their more effective displacement of expensive thermal generation due to the merit-order effect.
[…]
We further investigate the latter finding by analyzing the percentage relative difference between average real and counterfactual hourly quantity profiles, as stemming from the outcomes displayed in Figure 15, and by applying the smoothing procedure described at the beginning of this Section. Similarly to the hourly percentage impact on equilibrium quantities defined in Eq. (22), this indicator is computed as the relative deviation between average observed and synthetic quantities over the selected time window.
Formally, for each settlement hour $h$, let $\bar{Q}_{actual,h}$ and $\bar{Q}_{synt,h}$ denote the average cleared quantities in the actual and counterfactual markets, respectively, computed over all days in the reference period. The relative difference is then defined as:
$\text{RelDiff}_{h}=\frac{\bar{Q}_{actual,h}-\bar{Q}_{synt,h}}{\bar{Q}_{synt,h}}\times 100.$ (23)
By construction, a positive relative difference indicates that, on average, market equilibrium volumes are higher in the presence of RECs compared to the no-REC baseline, while a negative value signals a contraction in traded quantities, typically associated with a prevailing effect of REC self-consumption. This metric enables a transparent assessment of both the magnitude and the temporal (intra-day and seasonal) heterogeneity of REC-driven deviations from counterfactual market equilibria.
Yet, Figure 16 shows that both the HW (sc45.HW.2027) and the BU (sc45.BU.2027) scenarios display a positive relative percentage effect of RECs on actual market quantities. Conversely, the Policy scenario (sc45.2027) falls under the zero line for several settlement periods, thus suggesting that RECs have the potential to reduce traded market volumes in the Italian DA power market under large levels of deployment. Nevertheless, the magnitude of such effect has a strong seasonal dependency. Indeed, the largest magnitude of relative reduction in DA equilibrium quantities occurs in January during off-peak hours (10 a.m.–14 p.m.), as a result of the aforementioned ”REC winter effect”. These reductions during off-peak hours are related to loads from institutional REC members (public, SME, NPO).
Figure 16: Percentage relative difference of average hourly impact on quantities from RECs in the NORD. The outcomes are disentangled by settlement period, and displayed by month for each designed scenario.
[…]
Figure 17: Percentage relative difference between actual and counterfactual average hourly quantities in the NORD for each month, by hour. Scenario: sc45.BU.2027.
As for the Policy scenario (Figure 19 ), the impact of RECs in April would oscillate around the zero-line resembling ”zig-zags” until midday, thus witnessing a balanced effect between lower market demand (due to larger self-consumption by RECs) and higher electricity supply by RECs. Instead, the impact of RECs would be more pronounced during the rest of the day. Interestingly, the effect of RECs in July is the opposite, indicating a prevailing effect of energy injected by RECs until 1 p.m., to then switch back to the dominance of the effect of reduced market demand resembling ”zig-zags”, which could be explained by the strong need of power consumption for space cooling during summer. Indeed, we notice that the green line constantly falls before dropping below zero after 2 pm. We observe that the effect of reduced market demand (explicitly evident within the Policy scenario) mimics the hourly self-consumption rates of different prosumers’ categories. For example, in July, self-consumption rates in the afternoon (1pm - 7pm) for all prosumers’ categories are higher than the morning self-consumption rates (6am - 12pm). [^35] Public prosumers raise self-consumption rates on average at 2.3%, SME prosumers at 2.7%, residential prosumers at 18.5%, and NPO prosumers at 1.1%. In contrast, in April, self-consumption rates are, on average, higher in the morning than in the afternoon. We think that the specific composition of prosumers in RECs (Figure 9) can also affect the impact’s dynamics.
Figure 18: Percentage relative difference between actual and counterfactual average hourly quantities in the NORD for each month, by hour. Scenario: sc45.HW.2027.
Figure 19: Percentage relative difference between actual and counterfactual average hourly quantities in the NORD for each month, by hour. Scenario: sc45.2027.
[…]
Figure 20: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekdays. Scenario: sc45.BU.2027.
Figure 21: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekends. Scenario: sc45.BU.2027.
Figure 22: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekdays. Scenario: sc45.HW.2027.
Figure 23: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekends. Scenario: sc45.HW.2027.
Figure 24: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekdays. Scenario: sc45.2027.
Figure 25: Percentage relative difference between actual and counterfactual average hourly quantities in NORD for each month during weekends. Scenario: sc45.2027.
4.2.2 Sensitivity analysis: mixed scenario
[…]
In the scenario where RECs are widely deployed (Policy scenario), the system would recognize the role of RECs in reducing wholesale power demand through increased self-consumption. This would not only shield members of RECs from market price volatility, but also reduce the need for costly grid infrastructure investments while consumers’ autonomy would be enhanced.
Figure 26: Profile of average hourly quantities in NORD for both actual and counterfactual scenarios. Period: January 2027. Comparison between mixed scenario and policy scenario: sc_\_mix1.2027 vs. sc55.2027.
Figure 27: Hourly percentage impact on equilibrium quantities by RECs: comparison between mixed scenario and 3d policy scenario: sc_\_mix1.2027 vs sc55.2027 for NORD. Period: January 2027.
5 Discussion
First, our results show that the impact of RECs on power market is focused primarily in the NORD and CSUD market zones, with current negligible or absent effects in the other five physical market zones. This finding is closely related to the current distribution of REC projects, which are predominantly located in such two areas (see Fig. 10). Relatively slow and complex authorization and acceptance processes, institutional burdens, and limited technical support in other areas prevent broader deployment, confirming the need for targeted support policies to unlock REC potential across all regions. The similar finding has been reported by Zhu et al. (2025) and Musolino et al. (2023).
Second, the temporal dimension of REC impacts reveals strong seasonal and hourly patterns, driven by the interaction between solar production profiles and the assumed self-consumption rates. Our simulations for BU and HW scenarios indicate that certain early morning and evening hours - especially in spring and summer months - exhibit a relatively significant increase in market quantities, thereby highlighting the potential for RECs to better fulfill peak demand pressures in a decarbonizing system. Similarly, although including BESS, Riaz et al. (2019) modeled a summer week in the Australian wholesale market. They investigated the aggregate effect of a large number of prosumers on the load profile and found that increased prosumer participation with BESS flattens demand profiles, enhancing voltage stability and reducing the need for gas peaking plants. However, the exception in their study is the scenario with low demand and excess RES generation. It leads to increased aggregate demand due to battery charging, thus decreasing the stability margins of the system. Boccard and Goetz (2025) argue that excessive RES generation from REC prosumers entails increased power exchanges with the grid, although they do not specify in which direction.
[…]
Fourth, our results for the policy scenario indicate that RECs, by enabling higher self-consumption, could also reduce the need for costly investments in grid infrastructure. This finding aligns with the conclusions by Fuentes González et al. (2022), who modeled the effect of energy communities on transmission grid costs in Chile and found substantial system-wide savings. Backe et al. (2022) came to similar conclusions regarding the expansion of projected transmission capacity of the grid in six EU countries by 2060. Studies by Boccard and Goetz (2025) and Sarfarazi et. al. (2020) reveal that RECs could reduce pressure on the distribution grid too. Nonetheless, this positive effect is contingent upon the size of a PV system and composition of REC members in the former study and pricing strategies and community BESS operation in the latter study. While our study does not explicitly focus on optimization of grid costs, the observed market impacts also suggest that RECs may alleviate pressure on the grid.
Fifth, nuanced PV sizing will be essential to fully realize the potential of REC deployment not only to reduce stress on the grid but also wholesale market prices. Boccard and Goertz (2025) found that the different sizing of PV systems in RECs could induce opposite impacts on the distribution system. For example, RECs with homogeneous profiles of residential prosumers, which install PV systems that cover 25% - 50% of their load, reduce power exchanges with the grid, whereas those who install PV systems that cover 50%–100% of their load increase power exchanges with the grid due to the mismatch between energy injecting and self-consuming. The latter situation could create reverse flows and thus force additional costs for the DSOs that eventually will be shifted onto the non-members of RECs. In contrast, RECs with greater number of residential heterogeneous prosumers and RECs with public prosumers, which install PV systems that cover 25% of their load, increase power exchanges with the grid, while those who install PV systems that cover 50% - 100% of their load reduce power exchanges with the grid. This drastically opposite impacts on power exchanges hint at importance of modeling not only diverse PV system sizes but also diverse categories of prosumers, as we do. In our simulations, we used real-world averaged PV sizes of REC prosumers and producers in Italy (Figure 10) differentiated by categories and electricity zones.
6 Conclusions and Policy Implications
Our results show that REC deployment generates non-negligible effects in specific zones of the market (notably North and Central-South), while its overall system-wide impact remains quite moderate. Seasonal and hourly patterns confirm that self-consumption is a key driver of economic benefits from REC deployment at macro scale. Specifically, the estimated impacts on market volumes lie within a narrow range, from a minimum of -0.19% (January, Policy Scenario) to a maximum of +1.16% (April, BU Scenario). While both BU and HW scenarios yield consistently positive but limited net effects (up to +1.16% and +0.89%, respectively), the Policy scenario reveals mixed evidence: the impacts remain close to zero in most of the represented months, but turn slightly negative in January, suggesting that higher REC self-consumption can, under certain conditions, reduce the volumes available in the wholesale market. This observation becomes especially relevant if looking at the peak/off-peak dynamics. The reduction in equilibrium volumes during off-peak hours due to institutional prosumers within RECs (public, SMEs, NPOs) can reduce the need for RES curtailment while still having additional RES capacity deployed. On the other side, the increase in volumes during peak hours in spring and summer reduce the need for expensive thermal peaking plants. In contrast, when the increase in volumes occurs during off-peak hours, such as in scenarios with smaller capacity deployed (BU and HW), RECs adversely affect the market outcomes.
Indeed, by extending the analysis to seasonal and intra-day variations, our results confirm that the magnitude of REC impacts varies substantially across months and hours. Under the BU scenario, April emerges as the most responsive period, as market volumes with RECs surpass volumes without RECs by up to 3% during weekdays, while July exhibits a steadier impact close to 1% during peak hours. Conversely, cold months display more muted effects. The main factors behind this trend are the lower solar irradiation and the greater heating needs supplied by electricity during cold months. Lastly, when increasing the self-consumption rates for all categories of prosumers (from 45% to 55%) in line with Italian policy targets for RECs (sc55.2027), the results yield an overall average reduction of about -0.3% on the equilibrium volumes of the Italian DA market, which is on 0.11% greater than the average reduction of -0.19% in the 45% self-consumption scenario. This finding indicates that the increase in self-consumption rates for RECs deepens the reduction in equilibrium quantities.
Standalone systems are included in our REC modeling corresponding to the actual deployment status in Italy. The reducing effect of RECs’ self-consumption on zonal market quantities also suggests that both the volumes of energy injected into the grid from renewable standalone plants and volumes from prosumers’ plants within a REC could be offset by the self-consumption of REC prosumers. This finding contributes to the debate on the effects of the renewable-dominated system with high-electrified energy demand (incl. heating, cooling and transport) on wholesale market outcomes (Böttger and Härtel, 2022; Backe et al., 2022; Riaz et al., 2019).
Based on these insights, we outline several policy recommendations:
  1. 1.
    The greater the capacity deployed by RECs, the more pronounced the impact of institutional members (public entities, small and medium enterprises, and non-profit organizations), due to their traditionally higher volumes of self-consumed energy during off-peak hours compared with residential members (Figures 4, 5). The current Italian policy incentivizes smaller-capacity installations over larger ones by applying rigid capacity thresholds (see TIP in Figure 1 and Appendix 7.2), irrespectively of a prosumer category. However, rather than injecting energy into the grid during off-peak hours, rooftop installations that cover the loads of institutional members can greatly benefit the system. Therefore, the level of incentivization could be tailored to the extent of PV coverage of a building’s loads, instead of relying on rigid capacity thresholds. Such a measure could yield additional welfare gains as the electrification of public and industrial heating accelerates.
  2. 2.
    Institutional REC members inject more energy into the grid during peak hours in spring and summer than residential members, who tend to self-consume more energy during this period. Consequently, institutional REC members help meet peak demand in the system. In addition to public prosumers and non-profit organizations, small and medium-sized enterprises also exhibit load profiles that differ from those of residential prosumers. However, the Italian incentive scheme for shared energy does not make participation in RECs economically attractive for SMEs (see Koltunov et al. (2025)). The 50% reduction in the premium tariff for SMEs that benefit from capital subsidies constitutes one of the main obstacles. If such restrictive rules are relaxed, system-wide benefits from new RECs could increase.
  3. 3.
    According to the logic of Italian policy (MASE, 2023) and regulation (GSE S.p.A., 2025a; Schiavo et al., 2022), the economic systemic rationale behind RECs lies in the avoided use of the grid resulting from the simultaneous local injection and consumption of energy. We also demonstrate that reductions in market equilibrium volumes—particularly pronounced during the cold months—could help the system avoid RES curtailment when greater REC-driven capacity is deployed (i.e., 5 GW or more). This additional systemic rationale should be recognized by policy and thus valued/remunerated accordingly.
  4. 4.
    The injection of energy from both standalone facilities and rooftop solar installations can be offset by local self-consumption at the zonal scale. Since feed-in tariffs and feed-in premiums for large-scale renewables have been progressively phased out in the EU, the incentive for energy communities could also be used to support the rollout of standalone plants with capacities of up to 1 MW but in a more location-efficient and socially inclusive manner. Policymakers may therefore consider raising the maximum capacity threshold (i.e 1 MW) for individual REC configurations in order to enhance their economic viability for REC members as well as for third-party entities [^36] Participation in RECs is not open to large companies or companies whose primary activity belongs to energy sector. Therefore, utilities and ESCOs tend to participate in ”energy sharing” scheme as third-party entities. This means they cannot directly benefit from capital subsidy and from incentive on shared energy but can request lease payments if they own a renewable installation and/or charge configuration management fee to RECs.
  5. 5.
    By adopting battery storage within energy communities, even greater reductions in wholesale market equilibrium quantities can be achieved. The discharge of community-owned batteries to meet the needs of REC members generates substantial systemic benefits during peak hours. In this way, costly and polluting thermal generation can be displaced by clean and social energy. Still, as was discussed by many authors (Soini et al., 2020; Schick et al., 2022; Chen et al., 2023b; Sarfarazi et al., 2020), the specific modality of a storage system ultimately determines the scale and direction of this systemic impact. In Italy, capital subsidies are limited to 40% of the costs incurred by RECs, a design feature that discourages the installation of costly storage. To avoid additional public spending while still promoting systemic efficiency, policy could specifically incentivize integrated photovoltaic–battery systems in regions with excess renewable generation (e.g., Southern Italy).
  6. 6.
    RECs have historically been regarded as collective action initiatives created primarily by citizens and for citizens (e.g., energy cooperatives). Recently, Italian policy has also allowed institutional actors to participate in RECs. The load profiles of institutional prosumers enable greater systemic benefits from REC deployment—particularly when appropriately aggregated with residential loads—than those associated with multiple residential prosumers. Consequently, other countries may also consider permitting institutional actors to participate in RECs.
This paper is not exempt from limitations. First, the calibration of scenarios relies on publicly available data on REC uptake and profiles, standard self-consumption rates, and the announced policy targets by 2027. Second, behavioral aspects and demand-side management responses are not fully captured, although they may play a decisive role in the future and in dynamic adjustments in demand profiles, mirroring peculiar weather conditions and real-time modifications. Third, the projected impact of REC deployment is evaluated under the assumption that the capacities of other RES technologies remain fixed. This allows us to isolate the marginal contribution of the REC deployment to market equilibrium, although the actual system impacts could differ if broader RES additions were accounted for. Fourth, we disregard consumers who participate in RECs, although theoretically they might amend its consumption profiles to maximize the incentive for ”shared energy” when joining RECs. However, as stated in Section 1, this degree of coordination has not yet been implemented in the Italian energy communities, moreover, at a substantial scale.
Despite limitations, our methodology has a wide replication potential. The synthetic counterfactual approach can be adapted to other EU (and non-EU) electricity markets to assess the role of RECs (or other heterogeneous distributed energy resources) on day-ahead market equilibrium. This would allow for a broader comparative insight into how local RECs can reshape the aggregate electricity demand and supply, price signals, grid requirements and market resilience. A robustness check under an “accelerated battery adoption scenario” could further refine our estimates, offering a valuable sensitivity benchmark for future policy analysis. Similarly to Boccard and Goertz (2025), our future research could also adopt sensitivity tests based on differentiating PV sizes, which could suggest insights related to PV sizing for energy communities and its consequences for changes in a wholesale equilibrium.
Finally, our findings provide the timely inputs for energy and environmental policy-making. The diffusion of RECs could foster systemic efficiency, energy justice and independence, particularly in current times of geopolitical uncertainty and high energy costs, as highlighted during the 2022 energy crisis. However, their legacy will depend on complementary measures: efficient policy and regulation, harmonized governance frameworks, stronger support for storage and demand-side flexibility, and access to financing resources. The Italian government extended a deadline that grants €2.2 billion in subsidies for new installations within RECs to six additional months, with a possible extension (Camera dei deputati, 2025). This optimistic signal suggests that actual REC deployment in Italy by December 2027 will not only match our ”Half-Way” scenario but also make the 5 GW ”Policy scenario” more attainable.
Supplementary material
Our database of Italian RECs is available at the following GitHub link:
https://github.com/maksym-koltunov/energy-communities-Italy.git
Acknowledgements
The authors are deeply grateful to Prof. Alessandro Massi Pavan from the Department of Engineering and Architecture and to the Interdepartmental Center for Energy, Environment and Transport ”Giacomo Ciamician”, University of Trieste, for the great organizational and networking support during this study.
7 Appendices
7.1 Appendix A. Glossary
Acronym Description
BESS Battery Energy Storage System
DA Day-ahead
CO2 Carbon dioxide
DER Distributed Energy Resources
DSO Distribution System Operator
EU European Union
GW Gigawatt
GSE Gestore Servizi Energetici
IEM Internal Electricity Market Directive
MW Megawatt
MWh Megawatt per hour
LCOE Levelized Cost Of Energy
MOE Merit order effect
NGO Non-governmental Organisation
NPO Non-profit Organisation
PV Photovoltaic
REC Renewable Energy Community
RED-II Second Renewable Energy Directive
RES Renewable Energy Sources
SME Small and medium-sized enterprises
TSO Transmission System Operator
USA United States of America
JARSC Jointly acting renewable self-consumers
7.2 Appendix B. Shared energy and cashback calculation
In Italian regulation, shared energy refers to the minimum amount between the total energy injected into the public grid by all EC members within a given hour and the total energy withdrawn (consumed) by all EC members within the same hour.
The generalized formula for calculating shared energy is:
$E_{\mathrm{sh}}(h)=\min\left\{\sum_{i=1}^{n}E^{\mathrm{in}}_{i,h}\;;\;\sum_{i=1}^{n}E^{\mathrm{out}}_{i,h}\right\}$ (24)
where:
$E_{\mathrm{sh}}$:
the shared energy,
$E^{\mathrm{out}}_{i,h}$:
the energy withdrawn from the grid in a given hour $h$ by the $i$-th member,
$E^{\mathrm{in}}_{i,h}$:
the energy fed into the grid in a given hour $h$ by the $i$-th member.
The cashback logics is given by ARERA (2022) and Technical Rules from GSE (2025a). It can be described as:
$\mathrm{Cashback}=E_{\mathrm{inj}}\,P_{\mathrm{sell}}\;+\;E_{\mathrm{sh}}\,TIP\;+\;V_{\mathrm{agus}}\;-\;\mathrm{GSE\;costs}$ (25)
where:
$P_{\mathrm{sell}}$:
price for which GSE is purchasing renewable energy from eligible subjects. Regulated by ARERA 280/07 and equivalent to the Prezzo Minimo Garantito (PMG).
TIP :
premium tariff for shared energy.
$V_{\mathrm{agus}}$:
valorization of avoided grid usage due to self-consumption.
GSE costs:
administrative costs incurred by GSE related to cashback calculation and remuneration.
It is critical to consider how distinct components of the cashback are calculated. The premium tariff itself is not a single value; the basic premium tariff for an EC consists of fixed and variable components:
[…]
where
  • 80-70-60 are fixed components of a premium tariff that according to CER Decree stands at 60–80 €/MWh depending on the asset capacity (if $P<200$ equals 80; if $200<P<600$ equals 70; if $P>600$ equals 60). PV systems in central and northern Italy get an extra premium of 4 and 10 €/MWh respectively.
  • $\max(0;\,180-P_{z})$ is a variable component with $P_{z}$ being the hourly zonal price of electricity. In the event of zonal prices being higher than 180 €/MWh the variable component is counted as 0 €/MWh. This is made to prevent excessive premium during the periods of wholesale price spikes.
  • 120-110-100 are components that vary by installed capacity. If $P>600$ equals 100 €/MWh, if $200<P<600$ equals 110 €/MWh, if $P>600$ equals 120 €/MWh. In the event of extremely low wholesale price $P_{z}$ this parameter prohibits exceeding the premium tariff beyond 100–120 €/MWh.
The valorization of avoided grid usage due to self-consumption Vagus includes two components: the reduction in grid costs due to self-consumption and the reduction in costs related to decreased grid losses from self-consumption. JARSCs are compensated for both components, whereas RECs receive compensation only for the reduction in grid costs due to self-consumption. This distinction arises because RECs still rely on the distribution grid for energy sharing, thereby incurring grid losses. In contrast, JARSCs operate within a single condominium, bypassing the distribution grid. The Vagus for JARSCs and RECs can be calculated using the following formulas:
$\text{JARSC: }V_{agus}=T+L$ (27)
$\text{REC: }V_{agus}=T$ (28)
where
  • $T$ is reduction of grid costs due to self-consumption
  • $L$ is reduction of costs related to decreased grid losses due to self-consumption
Reduction of grid costs due to self-consumption (T) is calculated (Schiavo et al., 2022) by formula:
$T=E_{sh}\times\bigl(TRASe+\max(BTAUm)\bigr)$ (29)
where
  • TRASe is a sum of transmission tariff for LV household consumers,
  • BTAUm is a distribution tariff for LV commercial and industrial consumers
Reduction of costs related to decreased grid losses due to self-consumption ($L$) is calculated (Schiavo et al., 2022) by formula:
$L=\sum_{i,h}\bigl(E_{sh}\times c_{agl}\times P_{z}\bigr)$ (30)
where
  • $c_{agl}$ is a coefficient for avoided grid losses of the voltage level to which generator/s are connected (2.6% for generators in LV grid, 1.2% for generators in MV grid),
  • $P_{z}$ is the hourly zonal energy price of the day-ahead market,
  • $E_{sh}$ is the energy shared by the $i$-th member
By identifying all components of the cashback, we can calculate the actual economic benefit for EC members:
$\textit{Benefit}=E_{sc}(P_{buy})+[E_{inj}(P_{sell})+E_{sh}(P_{r})+V_{agus}-GSE_{costs}]$ (31)
where
  • $E_{sc}$ is an amount of self-consumed energy by common loads for JARSCs and by prosumers for EC,
  • $P_{buy}$ is a retail price that a member of EC pays to its retailer.
7.3 Appendix C
Table 3: Economic impact of RECs deployment on other market stakeholders
Specific impacts of RECs on market stakeholders Direction of impact Reference
On Generators
RECs deployment reduces the need for additional large-scale generation due to a merit-order effect (MOE) of renewables therefore reducing conventional producers’ surplus. Detriment Robinson & Guayo (2022), Backe et al. (2022)
RECs operating BESS could displace fossil fuel generators at a faster scale than large renewable generators without BESS due to MOE at peak demand periods, therefore reducing conventional producers’ surplus. Detriment (generators) / Benefit (system) Robinson & Guayo (2022)
RECs operating BESS reduce RES curtailment due to self-consumption at peak production congested times which allows to decrease supply of a distributed generation when grid needs it most (in turn, large renewable generators do not need to be curtailed thus preserving revenues). Benefit Backe et al. (2022)
On Retailers
Aggregator RECs operating BESS achieve cost reduction by directly participating in a wholesale market bypassing retailers thus shrinking their potential revenues. Detriment Faia et al. (2021)
When a REC acts as an intermediary between consumers and retailer, the latter lose ability to diversify risks by differentiating over tariffs to different consumer groups. Detriment Biggar & Hesamzadeh (2022)
Diminished sales from REC members elicit tariffs inflating for remaining customers, which in turn trigger latter to become more self-sufficient too. Situation leads to a financial “death spiral” of retailers. Detriment Parag & Sovacool (2016), Sarfarazi et al. (2020)
If a community energy storage is owned by a retailer, then the real-time pricing optimized based on behavior of REC members could yield profits for the retailer while not increasing costs for any type of REC members and delivering profits for flexible REC members. However, real time tariffs need implementation of EMS unavailable at scale in many countries. Benefit Sarfarazi et al. (2020)
On DSOs
When a REC acts as a united entity supplied only by a single external retailer, effect of a non-simultaneitya of contracted capacity is eliminated, which in turn reduces overall payment to DSOsb. Benefit Biggar & Hesamzadeh (2022), Robinson & Guayo (2022)
Electrotechnical criterion for defining perimeter of connection based on connection to same HV/MV substation (appr. 10000 PODs) and geographical criterion based on zip-code (appr. 1000 PODs) simplify DSO interaction with RECs. Benefit Del Pizzo et al. (2022)
Electrotechnical criterion of defining perimeter of connection based on connection to same MV/LV substation (appr. 70 PODs) and geographical criterion based on municipality belonging (appr. 10 to 850000 PODs) impedes DSO interaction with numerous or heterogeneous RECs. Detriment Del Pizzo et al. (2022)
Presence of RECs in some markets can cause operability issues and grid disruption due to a more complicated control and management schemes. Detriment Parag & Sovacool (2016)
Possibility to automatically detect and respond to actual and emerging grid problems through aggregated RECs (similar to VPPs), that may increase system’s resilience and decrease renewable energy oversupply concerns. Benefit Parag & Sovacool (2016)
On non-member consumers
DSOs can incur revenue losses due to decreased volumetric (decreased electricity purchasing by RECs) and/or fixed (non-simultaneity effect) network payments as well as possible subsidized exemption of RECs from the network payments. These revenue losses would typically be shifted onto non-member consumers. Detriment Biggar & Hesamzadeh (2022), Del Pizzo et al. (2022), Robinson & Guayo (2022), Sarfarazi et al. (2020)
Take-up of renewable energy supplied by RECs and all the subsequent MOE decrease energy prices, thereby non-members can greatly benefit. Benefit Biggar & Hesamzadeh (2022)
If RECs obtain implicit subsidy from a government, the costs are usually cross subsidized to non-members through retail bills. Detriment Robinson & Guayo (2022)
  • a
    Example: if 10 consumers each contract a 10-kW capacity-based network charge, the total contracted capacity would be 100 kW. However, due to non-simultaneity, regulation typically requires that the system only meets a combined demand of 75 kW, effectively overcharging consumers by 25 kW, which covers fixed network costs. When consumers form a REC, they contract only for 75 kW, benefiting from the non-simultaneity effect themselves and avoiding the 25-kW surcharge (Robinson and Del Guayo, 2022).
  • b
    This impact can backfire to non-member consumers because DSOs, as regulated monopolies, would typically shift the reduced revenues to non-member consumer bills.
7.4 Appendix D
Impact of prosumers on electricity systems
Ventosa et al. (2005) provide an overview of the model categories used for electricity markets analysis: optimization, equilibrium, and simulation models. We found several studies that analyze the systemic impact of prosumers. While two studies are theoretical review works that provide an extensive narrative of benefits and challenges (Robinson and Arcos-Vargas, 2023; Simshauser et al., 2023), the remaining studies rely on optimization models.
Simshauser et al. (2023) illustrate the case of Queensland state, Australia, which has the highest PV rooftop adoption rate in the world, while Robinson and Arcos-Vargas (2023) present both positive and negative effects of prosumer proliferation, focusing on Spain. Both studies outline many implications of prosumer proliferation, which often coincide with REC impacts. Some of the most pronounced implications include an adverse effect on conventional generators and challenges faced by DSOs due to a voltage rises (e.g., damage to customers’ electrical appliances). Another drawback is that retailers rapidly lose market share as prosumers penetration increases. Studies also point to an ambiguous impact on non-prosumers, who on one hand experience rising bills due to cross-subsidies and increased distribution charges, and on the other hand benefit from the decrease in the energy component of their retail bill as the result of the MOE and reductions in fuel costs. For instance, new PV installations in Spain resulted in a reduction of the wholesale electricity price of 0.01 euro for every 25 MWp installed, which, when aggregated across the overall market in 2021, led to user savings of more than 100,000 euros per year (Robinson and Arcos-Vargas, 2023). In Australia, another positive impact on non-prosumers was significantly lower installation costs due to the substantial growth of PV installation companies. In addition, DERs can lead to an increased need for ancillary services, which in turn reduces the price of these services due to greater market liquidity. Robinson and Arcos-Vargas (2023) argue that when prosumers and producers are aggregated (similar to RECs), costs of the distribution grid can drop because aggregated agents can provide flexibility services to the network — something that is almost impossible with disorganized individual installations. The proliferation of individual prosumers can decrease ohmic losses in the distribution grid, but only up to a certain level; beyond that, significant reverse flows occur, increasing losses again. Nevertheless, losses in such situations remain lower than before any DERs were deployed (2023, p. 135). An ambiguous impact, as underscored by Robinson and Arcos-Vargas (2023), occurs in terms of the security of supply when DERs do not utilize storage and are not aggregated. In this scenario, network costs can increase if a system is planned from an N-1 deterministic perspective; however, if it is planned from a probabilistic perspective, prosumers could improve security of supply even without storage or aggregation.
The first group of empirical studies on the systemic impact of prosumers attempts to quantify the associated challenges. For example, Schick and Hufendieck (Schick and Hufendiek, 2023) investigate the distributional spatial effect of the German feed-in-tariff during the period 2000-2021. Aggregated across Germany, the feed-in-tariff led to a cost shift of more than 500 million euros onto traditional consumer households. In 2021, maximization of self-consumption accounted for approximately half of this total effect. Tsybina et al. (2023) explore strategic behaviour of prosumers (exercising market power) and their response to the allocation of network losses—either to demand-side or the supply-side—as well as the impact of net metering policies. The authors determine that prosumers sell more electricity when losses are allocated to the demand base, whereas when losses are allocated to the supply base, prosumers sell less electricity. Another key observation is that lower wholesale equilibrium prices occur when network losses are allocated to the demand side due to two main factors. First, incorporating losses into the retail price (demand side) keeps selling prices higher and incentivizes prosumers to inject electricity into the grid. Second, higher retail prices encourage both consumers and prosumers to adjust their consumption patterns, leading to a more efficient use of energy, thereby reducing peak demand and the need for expensive peaking plants. Chen et al. (2023b) compare net-metering, net-billing, and benchmark policies [^37] Benchmark policy in this study assumes prosumers selling at a wholesale equilibrium price and buying at a retail price, therefore same tariffs are implied for prosumers as for other agents., examining their differential effects on various aspects of system welfare. The authors find that social surplus under the benchmark policy is significantly higher than under the other policies, making the benchmark policy the most welfare-enhancing. Under net metering, the transmission tariff would be 33% higher compared to the benchmark case, whereas the transmission tariff under net billing is closer to the benchmark case, making net billing the second-best solution. Another key finding is that wholesale social surplus (excluding prosumer surplus) deteriorates when prosumers saturate the node due to a greater number of consumers converting to prosumers—an indication of the “death spiral” effect. Chen et al. (2023a) refine the optimization model used in their previous study to investigate the impact of prosumers on transmission charges and social surpluses under the benchmark case alone. The authors assume different deployment levels of aggregated prosumers and analyse scenarios of perfect and imperfect competition, with the latter allowing prosumers to exhibit market power. The study reveals that wholesale prices decline under both scenarios due to increased renewable dispatch and reduced demand by prosumers. However, under imperfect competition — where prosumers strategically maximize their individual economic optimum — a significant increase in transmission charge occurs at all levels of deployment, particularly in scenarios with a high saturation of prosumers at nodes. This, in turn, reduces overall welfare.
The second group of empirical studies examines solutions to mitigate network costs shifting onto non-prosumers. Schick et al. (2021) demonstrate that network allocation schemes based on peak-coincident network capacity utilization can more effectively incentivize distribution network-oriented behaviour while ensuring a fairer distribution of financial burden between prosuming and non-prosuming households compared to volumetric network charges. A subsequent study by the same authors (2022) finds that higher self-consumption, when operated at least partially in a grid-beneficial manner (e.g., coupled with storage capable of providing flexibility service), can enhance RES integration and reduce CO2 emissions while avoiding cost shifting onto consumers. This finding suggests that dispersed prosumers could contribute more effectively to the grid if coordinated through a REC. On the other hand, when prosumers focus solely on maximizing their individual economic optimum— without considering system economic optimum (e.g., when storage operation is entirely inflexible)—RES integration could decrease, leading to a substantial rise in system costs and CO2 emissions (Schick et al., 2022). These findings align with those of Chen et al. (2023b).
The third group of studies explores future scenarios characterized by a high penetration of renewables, including DERs. Böttger and Härtel (2022) investigate hypothetical German power day-ahead market in 2050, assuming the deployment of carbon-neutral electricity/heat/transport systems. Importantly, their study considers the role of various novel demand-electrification technologies, which contribute to both supply and demand. The authors find that variable RES market values can be stabilised by power demand from diverse electrification applications, including flexible storage, power-to-gas, and power-to-heat (heat pumps). Consequently, a fully renewable future does not necessarily imply the “cannibalization effect” and highly volatile wholesale prices. Soini et al. (2020) investigate the impact of prosumers’ BESS on power supply costs in the Swiss electricity market for 2030, comparing it to the status quo in 2015. Their findings indicate that when BESS operation is optimized from a power system perspective—through time-of-use tariffs, grid charging, power exchange minimization, and households’ aggregation via RECs—substantial cost savings can be achieved. These savings primarily result from the reduced generation requirements and the substitution of pumped-hydro storage with more efficient BESS. Conversely, when fully independent households optimize their self-consumption, costs increase. This outcome is aligned with the findings of both Schick et al. (2022) and Chen et al. (2023b). Finally, Riaz et al. (2019) analyze the effect of large-scale prosumer aggregation (including BESS) on wholesale demand positions and load profiles, with a specific focus on loadability [^38] Loadability - maximum amount of a load that a system can support before it collapses. and voltage stability. Their study reveals that the increased prosumer-BESS participation smooths demand profiles, enhances loadability and voltage stability, and reduces gas power plants utilization—-thus lowering wholesale electricity prices. However, in scenarios of low demand and excess RES generation, these benefits do not occur. In contrast, a higher RES penetration without BESS leads to reverse flows and a reduction in reactive power support capability, ultimately lowering system stability margins.